### v1

# 2d Conformal Field Theories on Magic Triangle

Abstract
The magic triangle due to Cvitanović and Deligne–Gross is an extension of the Freudenthal–Tits magic square of semisimple Lie algebras. In this paper, we identify all 2d rational conformal field theories associated to the magic triangle. These include various Wess–Zumino–Witten (WZW) models, Virasoro minimal models, compact bosons and their non-diagonal modular invariants. At level one, we find a two-parameter family of modular linear differential equation of fourth order whose solutions produce the affine characters of all elements in the magic triangle. We find a universal coset relation for the whole triangle which generalizes the dual pairs with respect to $(E_{8})_{1}$ in the Cvitanović–Deligne exceptional series. This leads to the dimension and degeneracy of each primary field and also to five atomic models which constitute all theories in the triangle. At level two, we find a special row of the triangle – the subexceptional series has novel $N=1$ supersymmetry, and the Neveu–Schwarz/Ramond characters satisfy a one-parameter family of fermionic modular linear differential equations. Moreover, we find many new coset constructions involving WZW models at higher levels.
1 Introduction
The magic triangle of groups was discovered in both theoretical physics Cvitanovic (1981) and mathematics Deligne and Gross (2002). It may be viewed as a triangular extension of the Freudenthal–Tits magic square $\mathfrak{m}(\mathbb{A}_{1},\mathbb{A}_{2})$ of semisimple Lie algebras and Lie groups, where $\mathbb{A}$ ranges over the four normed division algebras: the real numbers, complex numbers, quaternions, and octonions. From another perspective, the magic triangle can be regarded as a two-dimensional generalization of the celebrated Cvitanović–Deligne exceptional series Cvitanovic (1981, 2008); Deligne (1996). The original magic triangle of Lie groups introduced by Cvitanović (Cvitanovic, 1981, Table 6) is reproduced in Table 1. The magic triangle of Deligne and Gross appears in (Deligne and Gross, 2002, Table 1).
The magic triangle is symmetric, and each row (or equivalently, each column) is referred to as an exceptional series. In Table 1, the bottom row, terminating at $E_{8}$, is known as the Cvitanović–Deligne exceptional series. The row immediately above it, ending at $E_{7}$, is called the subexceptional series. The third row from the bottom, ending at $E_{6}$, is known as the Severi exceptional series, while the fourth row from the bottom, terminating at $F_{4}$, is referred to as the Severi-section exceptional series.
$A_{1}$ $1/3$
$U_{1}$ $A_{2}$ $1/2$
$A_{1}$ $G_{2}$ $2/3$
$U_{1}^{2}$ $A_{1}^{3}$ $D_{4}$ 11
[…]
$1/3$ $1/2$ $2/3$ 11 $3/2$ 22 33 55 $\nu\backslash\mu$
$\mathbb{R}$ $\mathbb{C}$ $\mathbb{H}$ $\mathbb{O}$
Table 1: The magic triangle of Lie groups of Cvitanović Cvitanovic (1981) and Deligne–Gross Deligne and Gross (2002). The lower right $4\times 4$ square is the Freudenthal–Tits magic square, where $a=2(\nu-1)$ is the dimension of the division algebra. The last four rows from the bottom up are Cvitanović–Deligne, sub, Severi and Severi-section exceptional series respectively.
Each exceptional series exhibits distinctive and remarkable properties in representation theory. For instance, the Cvitanović–Deligne exceptional series displays striking uniformity in the tensor product decompositions of adjoint representations, and the dimensions of its irreducible representations admit expressions as rational functions of the dual Coxeter number Deligne (1996); Cohen and de Man (1996); Deligne and de Man (1996). From the perspective of Vogel’s theory of the universal Lie algebra, both the Cvitanović–Deligne series and the subexceptional series correspond to straight lines in Vogel’s projective plane Vogel (1999). Moreover, the four exceptional series are deeply connected with geometry Landsberg and Manivel (2002, 2004, 2001). In particular, the terminology Severi exceptional series and Severi-section exceptional series reflects their geometric relationship with Severi varieties Landsberg and Manivel (2004).
Several conceptual issues arise in the original magic triangle displayed in Table 1. First, despite its name, the original magic triangle of Lie groups does not form a complete triangular array, as certain entries along the slope are missing. These absent elements were identified as discrete groups in Deligne and Gross (2002). A second issue concerns the common practice in the physics literature of adjoining an additional entry “Lee–Yang model” at the beginning of the Cvitanović–Deligne exceptional series, which further obscures the triangular structure. A third question is whether the famous intermediate Lie algebra $E_{7+1/2}$ Landsberg and Manivel (2006), lying between $E_{7}$ and $E_{8}$, can be incorporated into the Cvitanović–Deligne exceptional series in a manner compatible with the magic triangle. In this work, we address these three issues in a unified framework and show that their resolution naturally requires an extension of the original diagram, resulting in a genuinely triangular structure.
Our approach is based on two-dimensional rational conformal field theories (RCFTs) associated with the magic triangle. This may be viewed as a natural generalization of the work of Mathur, Mukhi, and Sen on the 2d RCFTs associated with the Cvitanović–Deligne exceptional series Mathur et al. (1988, 1989). From this perspective, the resulting structure can be interpreted as an affine analogue of the original magic triangle. The construction is particularly natural at level one; the resulting extended triangle is displayed in Table 2. For each Lie group appearing in the original magic triangle in Table 1, the corresponding entry in Table 2 is given by the associated Wess–Zumino–Witten model mostly at level one. The remaining entries consist of various minimal models, compact boson theories, and their non-diagonal modular invariants, which will be explained in detail below. In particular, we introduce an additional row ending at $E_{7+1/2}$, referred to as the intermediate exceptional series, which has been studied recently in Lee et al. (2024b).
0 1/5
0 $\mathrm{Vir}_{5,2}^{\rm eff}$ 1/4
0 $\mathrm{Vir}_{5,3}^{\rm eff}$ $(A_{1})_{1}$ 1/3
0 $(U_{1})_{3}$ $(U\!A_{1\frac{1}{2}})_{1}$ $(A_{2})_{1}$ 1/2
0 $\mathrm{Vir}_{6,5}^{\rm e}$ $(A_{1})_{3}$ $(AG_{1\frac{1}{2}})_{1}$ $(G_{2})_{1}$ 2/3
0 $D_{2\rm A}$ $M_{2}$ $(A_{1}^{\oplus 3})_{1}$ $(AD_{3\frac{1}{2}})_{1}$ $(D_{4})_{1}$ 1
0 $D_{2\rm A}$ $(A_{1})_{8}^{\rm e}$ $(A_{2})_{2}$ $(C_{3})_{1}$ $(C_{3\frac{1}{2}})_{1}$ $(F_{4})_{1}$ 3/2
0 $\mathrm{Vir}_{6,5}^{\rm e}$ $M_{2}$ $(A_{2})_{2}$ $(A_{2}^{\oplus 2})_{1}$ $(A_{5})_{1}$ $(A_{5\frac{1}{2}})_{1}$ $(E_{6})_{1}$ 2
0 $(U_{1})_{3}$ $(A_{1})_{3}$ $(A_{1}^{\oplus 3})_{1}$ $(C_{3})_{1}$ $(A_{5})_{1}$ $(D_{6})_{1}$ $(D_{6\frac{1}{2}})_{1}$ $(E_{7})_{1}$ 3
0 $\mathrm{Vir}_{5,3}^{\rm eff}$ $(U\!A_{1\frac{1}{2}})_{1}$ $(AG_{1\frac{1}{2}})_{1}$ $(AD_{3\frac{1}{2}})_{1}$ $(C_{3\frac{1}{2}})_{1}$ $(A_{5\frac{1}{2}})_{1}$ $(D_{6\frac{1}{2}})_{1}$ $(D_{6\frac{1}{2}\!\frac{1}{2}})_{1}$ $(E_{7\frac{1}{2}})_{1}$ 44
0 $\mathrm{Vir}_{5,2}^{\rm eff}$ $(A_{1})_{1}$ $(A_{2})_{1}$ $(G_{2})_{1}$ $(D_{4})_{1}$ $(F_{4})_{1}$ $(E_{6})_{1}$ $(E_{7})_{1}$ $(E_{7\frac{1}{2}})_{1}$ $(E_{8})_{1}$ 5
$1/5$ $1/4$ $1/3$ $1/2$ $2/3$ 11 $3/2$ 2 3 4 5 $\nu\backslash\mu$
Table 2: 2d CFTs associated to the extended magic triangle at level one.
The main results of this work can be summarized as follows.
1. We determine the 2d RCFT associated with each entry of the extended magic triangle at level one. The characters of these RCFTs satisfy a uniform two-parameter family of fourth-order holomorphic modular linear differential equations (MLDEs), given in (41).
2. The magic coset (48) provides a natural generalization of the dual pair structure with respect to $(E_{8})_{1}$ in the Cvitanović–Deligne exceptional series to the entire magic triangle.
3. The dimensions, degeneracies, and modular $S$-matrices of all theories in the magic triangle are given uniformly.
4. We show that all thirty theories on the level-one magic triangle can be decomposed into five elementary building blocks ${\rm Vir}^{\rm eff}_{5,2}$, ${\rm Vir}^{\rm eff}_{5,3}$, $(U_{1})_{3}$, ${\rm Vir}^{\rm e}_{6,5}$ and $D_{\rm 2A}$.
5. The subexceptional series at level two exhibits $\mathcal{N}=1$ supersymmetry. The corresponding Neveu–Schwarz sector characters satisfy a one-parameter family of fourth-order fermionic MLDEs.
6. Uniform coset constructions for $(\mathfrak{g}_{\rm CD})_{2}/{(\mathfrak{g}_{\rm sub})_{2}\times(A_{1})_{2}}$ and $(\mathfrak{g}_{\rm sub})_{2}/{(\mathfrak{g}_{\rm Severi})_{2}\times(U_{1})_{6}}$.
This paper is organized as follows. In Section 2, we review background material on exceptional series, the magic square, and the magic triangle, and provide a brief introduction to characters of 2d RCFTs and modular linear differential equations (MLDEs). In particular, we recall the seminal result of Mathur, Mukhi, and Sen Mathur et al. (1988) that the CFT-type solutions of second-order holomorphic MLDEs are precisely the level-one affine characters of the Cvitanović–Deligne exceptional series. In Section 3, we extend the framework of Mathur–Mukhi–Sen to the full magic triangle and investigate several structural properties of this generalization, which may be regarded as an affine realization of the magic triangle at level one. Section 4 is devoted to the magic triangle at level two, with particular emphasis on the exceptional series appearing in the lower rows which exhibits more symmetries. In Section 5, we discuss the magic triangle at arbitrary higher levels. In the end, we give some concluding remarks. In Appendix A, we explicitly present the data for the extended magic triangle and CFTs at level one. In Appendix B, we explicitly give the S-matrices for the CFTs at level one of CD exceptional series. In Appendix C, we discuss the spurious solutions of the 4th order MLDE for CD exceptional series.
We use $()^{\rm e}$ for a non-diagonal modular invariant of RCFT, i.e. a VOA extension. For example, ${\rm Vir}^{\rm e}_{6,5}$ is a non-diagonal modular invariant of Virasoro minimal model ${\rm Vir}_{6,5}$ that describes the critical three-state Potts model. We denote the RCFT associated to the triangle element $(\mu,\nu)$ at level $k$ as $\mathcal{T}^{(\mu,\nu)}_{k}$. Especially we denote the level-one theory as $\mathcal{T}^{(\mu,\nu)}$ for simplicity. We use subscript to denote the degeneracy of primary fields. We use $\rm eff$ to denote the effective theory of non-unitary minimal models.
2 Review on the magic triangle
2.1 Exceptional series and magic triangle
We begin by introducing a key parameter $\nu$ associated with the magic triangle and the exceptional series. Throughout this paper, $\nu$ takes values in the finite set
$\{1/5,1/4,1/3,1/2,2/3,1,3/2,2,3,4,5\}.$ (1)
They correspond to the following extended Cvitanović–Deligne exceptional series respectively
$e\subset A_{1/2}\subset A_{1}\subset A_{2}\subset G_{2}\subset D_{4}\subset F_{4}\subset E_{6}\subset E_{7}\subset E_{7+1/2}\subset E_{8}.$ (2)
Here $e$ is a trivial element. For the simple Lie algebras $\mathfrak{g}$ in the series, which constitute the original Cvitanović–Deligne exceptional series, the dimension and the dual Coxeter number Cvitanovic (1981, 2008); Deligne (1996) are given by $\nu$ as
$d_{\mathfrak{g}}$ $=\frac{2(5\nu-1)(6\nu+1)}{(\nu+1)},$ (3)
$h^{\vee}_{\mathfrak{g}}$ $=6\nu.$ (4)
Moreover, infinitely many representations of $\mathfrak{g}$ admit dimension formulas that can be expressed as rational functions of $\nu$ Cohen and de Man (1996); Landsberg and Manivel (2006).
The original CD exceptional series is extended to include intermediate Lie algebras $A_{1/2}$ and $E_{7+1/2}$, whose dual Coxeter number and dimension obey the same formula. To be specific, $A_{1/2}$ has dual Coxeter number $3/2$ and dimension $1$. In many senses, one can regard $A_{1/2}$ as Lie superalgebra $osp(1|2)$ which has dual Coxeter number $3/2$ and superdimension $1$. Then affine characters of $(A_{1/2})_{k}$ are just the supercharacters of $osp(1|2)_{k}$. An alternative interpretation is to regard $A_{1/2}$ as a formal object, denoted $LY$, such that the associated theory $(LY)k$ realizes the Lee–Yang model at level $k$, namely the effective Virasoro minimal model $\mathrm{Vir}^{\rm eff}_{(2k+3,2)}$ Duan et al. (2022). This correspondence follows from the fact that the supercharacters of $\mathfrak{osp}(1|2)_{k}$ agree, as $q$-series, with the characters of $\mathrm{Vir}^{\rm eff}_{(2k+3,2)}$. The $E_{7+1/2}$ is a famous intermediate Lie algebra with dual Coxeter number $24$ and dimension $190$. Its existence was noted both in physics Mathur et al. (1988); Cvitanovic (2008) and mathematics Cohen and de Man (1996), and it was subsequently constructed by Landsberg and Manivel using sextonions Landsberg and Manivel (2006).
The magic triangle is a two-dimensional symmetric generalization of the Cvitanović–Deligne exceptional series, parameterized by a pair $(\mu,\nu)$ with $\mu\nu\geq 1$. In the case $\mu\nu=1$, all entries reduce to trivial elements. One systematic construction of the magic triangle proceeds by placing the Cvitanovi’c–Deligne exceptional series along the bottom row and the rightmost column, and then defining each remaining entry as the centralizer of the corresponding row element inside the column element Deligne and Gross (2002). The original magic triangle introduced in (Cvitanovic, 1981, Table 6) is reproduced in Table 1. The subarray with $\mu,\nu>1$, highlighted in Table 1, is known as the Freudenthal–Tits magic square. The Lie algebras appearing in the magic square admit a uniform construction in terms of pairs of normed division algebras, $\mathfrak{m}(\mathbb{A}_{1},\mathbb{A}_{2})$ for $\mathbb{A}=\mathbb{R},\mathbb{C},\mathbb{H},\mathbb{O}$. This construction realizes the idea that all five exceptional Lie algebras are related to the octonions; see Baez (2002) for a comprehensive review. Upon including the sextonion algebra $\mathbb{S}$, the magic square can be further extended to incorporate several intermediate Lie algebras; we refer to Lee et al. (2024b) for a detailed discussion. In this context, it is sometimes convenient to replace the parameter $\nu$ by $a=2(\nu-1)$, which takes values in the set $\{-8/5,-3/2,-4/3,-1,-2/3,0,1,2,4,6,8\}.$ Here the numbers $1,2,4,6,8$ are the real dimension of $\mathbb{R},\mathbb{C},\mathbb{H},\mathbb{S},\mathbb{O}$. Finally, it is worth noting that the magic square has also appeared in the study of supergravity theories, see for example Anastasiou et al. (2014).
For the element ${\mathfrak{g}}^{(\mu,\nu)}$ in the original magic triangle in Table 1, the dimension and dual Coxeter number have the following universal formulas
$d_{\mathfrak{g}}^{(\mu,\nu)}$ $=\frac{12(\mu\nu-1)(\mu\nu+\mu+\nu-4)}{(\mu+1)(\nu+1)},$ (5)
$h^{\vee}_{\mathfrak{g}}{}^{(\mu,\nu)}$ $=\mu\nu+\mu+\nu-5.$ (6)
The dimension formula coincides with that of the magic square derived in Landsberg and Manivel (2006). In particular, by specializing to $\mu=5$, one recovers the expressions for the dimensions and dual Coxeter numbers of the Cvitanović–Deligne exceptional series. The original magic triangle of Cvitanović and of Deligne–Gross does not include the cases $\mu,\nu=1/4$ or $4$. Incorporating these values leads naturally to an extended magic triangle. Remarkably, the universal formulas (5) and (6) continue to hold throughout this extension. For $\nu=1/4$, the extension introduces a single additional element, namely $A_{1/2}$. For $\nu=4$, one obtains the intermediate exceptional series, whose dimensions and dual Coxeter numbers were determined in Lee et al. (2024b) and are consistent with the general formulas (5) and (6). The dual Coxeter numbers and dimensions of all entries in the extended magic triangle are summarized in Table 5.
Several entries of the extended magic triangle require additional clarification. The formal groups ${\mathfrak{g}}^{(3/2,1)}$ and ${\mathfrak{g}}^{(2,2/3)}$ both have vanishing dimension and dual Coxeter number equal to 1-1. In Deligne and Gross (2002), these entries are realized concretely by the finite groups $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ and $\mathbb{Z}_{3}$, respectively. The formal group ${\mathfrak{g}}^{(4,1/3)}$ has dimension $1$ and dual Coxeter number $2/3$. It may be interpreted as a formal object, denoted $IM$, such that the associated theory $(IM)_{k}$ realizes the effective minimal model $\mathrm{Vir}^{\rm eff}_{(3k+2,3)}$ Lee et al. (2024b). There also exist three entries whose underlying Lie algebras have standard dimensions but nonstandard dual Coxeter numbers. These are ${\mathfrak{g}}^{(3,2/3)}=A_{1}$, ${\mathfrak{g}}^{(3/2,3/2)}=A_{1}$ and ${\mathfrak{g}}^{(2,3/2)}=A_{2}$. According to the general formula (6), their dual Coxeter numbers are given by $2/3$, $1/4$, and $3/2$, respectively. Equivalently, the invariant bilinear forms on these Lie algebras are rescaled by factors of $3$, 88, and $2$. Consequently, when constructing the associated Wess–Zumino–Witten models, the levels must be rescaled by the same factors, namely $3$, 88, and $2$, respectively.
2.2 Modular linear differential equation
In this section, we briefly review modular linear differential equations (MLDEs) and their fermionic version, namely MLDEs associated with congruence subgroups of level two. There exist several equivalent formulations of MLDEs, including those based on ordinary derivatives, Serre derivatives, and Rankin–Cohen brackets; see, for example, Nagatomo et al. (2024). For our purposes, it is most convenient to work with the Serre derivative
$D_{k}=q\frac{d}{dq}-\frac{k}{12}E_{2},$ (7)
which maps a holomorphic modular form of weight $k$ to a holomorphic modular form of weight $k+2$. A modular linear differential equation of degree $d$ can be written in the form
$\phi_{2l}D^{d}+\phi_{2l+2}D^{d-1}+\phi_{2l+4}D^{d-2}+\dots+\phi_{2l+2d}]\chi=0,$ (8)
where the coefficients $\phi_{k}$ are holomorphic modular forms of weight $k$ for $SL(2,\mathbb{Z})$, generated by the Eisenstein series $E_{4}$ and $E_{6}$. The non-negative integer $l$ is referred to as the index of the MLDE. When $l=0$, the equation is called holomorphic or monic; otherwise it is meromorphic.
We are interested in the special situation where $\chi$ is a modular form of weight zero on certain congruence subgroup of $SL(2,\mathbb{Z})$. In particular, we focus on solutions of the form
$q^{\alpha}(a_{0}+a_{1}q+a_{2}q^{2}+\dots),$
where $\alpha$ is a rational number, and all Fourier coefficients $a_{i}$ are non-negative integers. Such solutions are referred to as CFT-type solutions of MLDEs and constitute a necessary condition for $\chi$ to be admissible as a character of a two-dimensional rational conformal field theory (RCFT). The modular properties of RCFT characters play a vital role in the study of two-dimensional conformal field theories. The full set of characters of a 2d RCFT forms a vector-valued modular form for $SL(2,\mathbb{Z})$ Zhu (1996), and the congruence subgroup property of RCFT characters was established in Dong et al. (2015). The systematic study of MLDEs satisfied by RCFT characters was initiated by Eguchi and Ooguri Eguchi and Ooguri (1988) and further developed in Anderson and Moore (1988); Mathur et al. (1988).
Suppose that the MLDE (8) admits $d$ linearly independent CFT-type solutions forming a vector-valued modular form. The Wronskian analysis of Mathur et al. (1988) implies that the index $l$ is related to the degree $d$ and the exponents $\alpha_{i}$ by
$\frac{l}{6}=\frac{d(d-1)}{12}-\sum_{i=0}^{d-1}\alpha_{i}.$ (9)
When these solutions are identified with the characters of a 2d RCFT, the exponents are related to the central charge $c$ and conformal weights $h_{i}$ by
$\alpha_{i}=-\frac{c}{24}+h_{i}.$
One of the important applications of MLDE is to classify 2d RCFTs with given number of characters, or bootstrap new 2d RCFTs. This program requires a classification of CFT-type solutions for given pairs $(d,l)$. In the next subsection, we review the classical work of Mathur, Mukhi, and Sen on MLDEs with $(d,l)=(2,0)$ Mathur et al. (1988). More recent progress on the classification of CFT-type MLDE solutions can be found, for example, in Kaidi et al. (2021).
It is natural to generalize the properties of the characters of 2d RCFTs to the case of 2d fermionic RCFTs. In the presence of fermions, one must distinguish between periodic (Ramond) and anti-periodic (Neveu–Schwarz) boundary conditions along a circle, leading to four sectors depending on the boundary conditions along the two cycles of the torus. Accordingly, fermionic RCFTs admit four types of characters: NS, $\widetilde{\rm NS}$, R, and $\widetilde{\rm R}$. These transform as vector-valued modular forms for the congruence subgroups $\Gamma_{\theta}(2)$, $\Gamma^{0}(2)$, $\Gamma_{0}(2)$, and $SL(2,\mathbb{Z})$, respectively; see, for instance, Bae et al. (2021a); Duan et al. (2023). The fermionic characters of a fermionic RCFT satisfy corresponding fermionic MLDEs. For example, the Neveu–Schwarz characters obey an MLDE of the form (8), where the coefficients $\phi_{k}$ are holomorphic modular forms of weight $k$ for the congruence subgroup $\Gamma_{\theta}(2)$.
2.3 Mathur–Mukhi–Sen’s discovery
The simplest nontrivial holomorphic modular linear differential equations are of second order. In Mathur et al. (1988, 1989), Mathur, Mukhi, and Sen classified all CFT-type solutions of the second-order holomorphic MLDE
$D^{2}+\lambda E_{4}]\chi=0.$ (10)
This equation is also known in the literature as the Kaneko–Zagier equation Kaneko and Zagier (1998). Mathur, Mukhi, and Sen showed that only ten discrete values of the parameter $\lambda$ admit CFT-type solutions. Moreover, for each such value of $\lambda$, the corresponding solutions coincide precisely with the characters of level-one Wess–Zumino–Witten models associated with the extended Cvitanović–Deligne exceptional series:
$\mathrm{Vir}_{5,2}^{\rm eff},\ (A_{1})_{1},\ (A_{2})_{1},\ (G_{2})_{1},\ (D_{4})_{1},\ (F_{4})_{1},\ (E_{6})_{1},\ (E_{7})_{1},\ (E_{7+1/2})_{1},\ (E_{8})_{1}.$ (11)
We refer to the sequence (11) as the Mathur–Mukhi–Sen series, and denote them as $\mathcal{T}^{(\nu,5)}$ or $\mathcal{T}^{(5,\nu)}$ theories with $\nu$ defined earlier associated to the CD exceptional series. This notation will be consistent with magic triangle at level one in Table 2. The associated CFT data are summarized in Table 3. Here $\mathrm{Vir}_{5,2}^{\rm eff}$ denotes the effective Lee–Yang model with effective central charge $2/5$. The entry $(E_{7+1/2})_{1}$ corresponds to the solution of the MLDE with central charge $c=38/5$. Mathur, Mukhi, and Sen observed that, if interpreted as a level-one WZW model, the corresponding symmetry algebra would have dimension $190$ and dual Coxeter number $24$, and hence could not arise from a semisimple Lie algebra. It was subsequently shown that these solutions can be realized as the characters of an intermediate vertex subalgebra of $(E_{8})_{1}$ Kawasetsu (2014). An alternative interpretation identifies them as the characters of a minimal $\mathcal{W}$-algebra with central charge $32/5$ Kawasetsu (2018). The $(E_{8})_{1}$ theory has the single vacuum character $j(\tau)^{1/3}$ Mathur et al. (1988). In this case, the second order MLDE has a spurious solution, see the Appendix C.
$\nu$ $1/4$ $1/3$ $1/2$ $2/3$ 11 $3/2$ 2 3 4 5
$\mathrm{Vir}_{5,2}^{\rm eff}$ $(A_{1})_{1}$ $(A_{2})_{1}$ $(G_{2})_{1}$ $(D_{4})_{1}$ $(F_{4})_{1}$ $(E_{6})_{1}$ $(E_{7})_{1}$ $(E_{7\frac{1}{2}})_{1}$ $(E_{8})_{1}$
$c$ $\frac{2}{5}$ 1 2 $\frac{14}{5}$ 4 $\frac{26}{5}$ 6 7 $\frac{38}{5}$ 8
$h$ $\frac{1}{5}$ $\frac{1}{4}$ $\frac{1}{3}$ $\frac{2}{5}$ $\frac{1}{2}$ $\frac{3}{5}$ $\frac{2}{3}$ $\frac{3}{4}$ $\frac{4}{5}$ -
$\mathcal{N}$ 1 1 2 1 3 1 2 1 1 -
$R$ 1 2 3 7 8 26 27 56 57 -
Table 3: 2d CFTs associated to the extended CD exceptional series at level 11.
Recall that the Wess–Zumino–Witten model $(\mathfrak{g})k$ has central charge
$c_{k}=\frac{kd_{\mathfrak{g}}}{k+h^{\vee}}.$ (12)
For the extended Cvitanovi’c–Deligne exceptional series, both the dimension $d_{\mathfrak{g}}$ and the dual Coxeter number $h^{\vee}$ can be expressed as functions of the parameter $\nu$. At level one, the corresponding central charge and the conformal weight of the non-vacuum primary field are given by
$c=\frac{2(5\nu-1)}{(\nu+1)},\qquad h=\frac{\nu}{\nu+1}.$ (13)
In terms of the parameter $\nu$, the associated second-order holomorphic MLDE can be written compactly as $\Delta_{\nu}\chi=0$, where $\Delta_{\nu}\chi=0$, where
$\Delta_{\nu}=D^{2}-\frac{(5\nu-1)(7\nu+1)}{144(\nu+1)^{2}}E_{4}.$ (14)
The modular invariant partition function of the 2d RCFT on torus takes the form
$Z=|\chi_{0}|^{2}+{\mathcal{N}}|\chi_{1}|^{2},$ (15)
where $\mathcal{N}$ is natural integer, capturing the degeneracy of the non-vaccum primary operator. The characters of the Mathur–Mukhi–Sen series admit the expansions
$\chi_{0}$ $=q^{-c/24}(1+d_{\mathfrak{g}}q+\cdots),$
$\chi_{1}$ $=q^{-c/24+h}(R+\cdots),$ (16)
where $R$ denotes the dimension of the fundamental representation of $\mathfrak{g}$. The $S$-matrix arises from the modular transformation of the characters,
$\left(\begin{array}[]{c}\chi_{0}\\ \chi_{1}\end{array}\right)(-1/\tau)=S\cdot\left(\begin{array}[]{c}\chi_{0}\\ \chi_{1}\end{array}\right)(\tau).$ (21)
As a $2\times 2$ matrix, the $S$-matrix satisfies $\det S_{(\nu)}=-1$ and
$S_{(\nu)}^{T}\cdot\left(\begin{array}[]{cc}1&0\\ 0&\mathcal{N}_{(\nu)}\end{array}\right)\cdot S_{(\nu)}=\left(\begin{array}[]{cc}1&0\\ 0&\mathcal{N}_{(\nu)}\end{array}\right).$ (26)
The explicit form of the $S$-matrices for the Mathur–Mukhi–Sen series is presented in Appendix B. More generally, one may consider the full modular $S_{(\nu)}$-matrix as a $(1+\mathcal{N}_{(\nu)})$-dimensional unitary matrix.
It is well known that for the CD exceptional series at level $1$, the RCFTs pair together as a coset with respect to $\mathcal{T}^{(5,5)}=(E_{8})_{1}$ Mathur et al. (1988). These cosets are
$(E_{7})_{1}=\frac{(E_{8})_{1}}{(A_{1})_{1}},\quad(E_{6})_{1}=\frac{(E_{8})_{1}}{(A_{2})_{1}},\quad(F_{4})_{1}=\frac{(E_{8})_{1}}{(G_{2})_{1}},\quad(D_{4})_{1}=\frac{(E_{8})_{1}}{(D_{4})_{1}},\quad$ (27)
In our notation, these cosets can be written uniformly as
$\mathcal{T}^{(5,\nu)}=\frac{(E_{8})_{1}}{\mathcal{T}^{(1/\nu,5)}}.$ (28)
The non-vacuum conformal weight $h_{\nu}$ of $\mathcal{T}_{5,\nu}$ satisfies the following relation $h_{\nu}+h_{1/\nu}=1.$ The degeneracy $\mathcal{N}_{(\nu)}$ and $S$-matrix $S_{(\nu)}$ satisfy
$\mathcal{N}_{(\nu)}=\mathcal{N}_{(1/\nu)},\ \ S_{(\nu)}=S_{(1/\nu)}.$ (29)
The characters satisfy the relation
$\chi_{0}^{(\nu)}\chi_{0}^{(1/\nu)}+\mathcal{N}_{(\nu)}\chi_{1}^{(\nu)}\chi_{1}^{(1/\nu)}=j^{1/3},$ (30)
which is $S$-invariant. The above bilinear relation in the subleading $q$-order implies
$\mathcal{N}_{(\nu)}R_{(\nu)}R_{(1/\nu)}=312-60(\nu+1/\nu).$ (31)
From the study of the solution in terms of the hypergeometric function Naculich (1989) with $h=\nu/(\nu+1)$ being the conformal weight of the non-vacuum primary operator, the explicit formula for $\mathcal{N}_{(\nu)}R_{(\nu)}^{2}$ is known as
$\mathcal{N}_{(\nu)}R_{(\nu)}^{2}=2^{16h}(4\sin^{2}(\pi h)-1)\bigg[\frac{\Gamma(1-2h)\Gamma(1/2+3h)}{\Gamma(1+2h)\Gamma(1/2-h)}\bigg]^{2}.$ (32)
Multiplying $\mathcal{N}_{(\nu)}R_{(1/\nu)}^{2}$ one can also obtain the relation (31).
The characters of WZW models for CD exceptional series at level one have many other interesting properties such as Hecke relations Harvey and Wu (2018). Besides, it is recently shown that the vacuum character of WZW models for CD exceptional series at level 1 can be constructed from the monodromy trace of some Argyres–Douglas theories Kim and Song (2024).
3 Magic triangle at level one
A principal motivation of this work is to extend the discovery of Mathur, Mukhi, and Sen concerning second-order holomorphic MLDEs and the Cvitanovi’c–Deligne exceptional series to the full magic triangle. To this end, we begin by analyzing the level-one Wess–Zumino–Witten models associated with the ordinary Lie groups appearing in the original magic triangle of Table 1. For each resulting rational conformal field theory, we compute the characters and show that they satisfy a uniform two-parameter family of fourth-order holomorphic MLDEs.
Assuming that this MLDE framework extends to the enlarged magic triangle, we solve for the characters corresponding to the previously unknown entries and identify the associated 2d RCFTs. The complete set of resulting theories is summarized in Table 2, which we interpret as the extended magic triangle at level one. We further uncover several structural properties of this level-one triangle, including the magic coset, magic bilinear relation and matrix-valued modularity of the characters.
3.1 The fourth order MLDE
Let us first consider the level-one theories for the original magic triangle in Table 1. Substituting the dimension and dual Coxeter number formulas (5), (6) into the central charge formula of WZW model (12), we obtain the central charge for the level-one theory $\mathcal{T}^{(\mu,\nu)}$ as
$c^{(\mu,\nu)}=\frac{12(\mu\nu-1)}{(\mu+1)(\nu+1)}.$ (33)
We observe that each theory $\mathcal{T}^{(\mu,\nu)}$ admits at most four distinct characters. With the exception of the Cvitanović–Deligne exceptional series, the theory $\mathcal{T}^{(\mu,\nu)}$ with $\mu\neq\nu$ generically possesses four inequivalent characters. Their conformal weights take the universal form
$h_{i}^{(\mu,\nu)}=0,\ \frac{\mu\nu-1}{(\mu+1)(\nu+1)},\ \frac{\nu}{\nu+1},\ \frac{\mu}{\mu+1}.$ (34)
The exponents $\alpha_{i}=-c/24+h_{i}$ in the corresponding characters $\chi_{i}$ take values,
$\alpha_{0}=-\alpha_{1}=\frac{1-\mu\nu}{2(\mu+1)(\nu+1)},\quad\alpha_{2}=\frac{\mu\nu+2\nu+1}{2(\mu+1)(\nu+1)},\quad\alpha_{3}=\frac{\mu\nu+2\mu+1}{2(\mu+1)(\nu+1)}.$ (35)
Denote the characters of the theory $\mathcal{T}^{(\mu,\nu)}$ as
$\chi_{0}$ $=q^{\alpha_{0}}\Big(1+\sum_{n=1}^{\infty}a_{n}q^{n}\Big)\ \ {\rm with}\ a_{1}=d_{\mathfrak{g}},$ (36)
$\chi_{1}$ $=q^{\alpha_{1}}\Big(R_{1}+\sum_{n=1}^{\infty}b_{n}q^{n}\Big),$ (37)
$\chi_{2}$ $=q^{\alpha_{2}}\Big(R_{2}+\sum_{n=1}^{\infty}c_{n}q^{n}\Big),$ (38)
$\chi_{3}$ $=q^{\alpha_{3}}\Big(R_{3}+\sum_{n=1}^{\infty}d_{n}q^{n}\Big),$ (39)
where $R_{i}$ are the dimensions of the irreducible representation of $\mathfrak{g}^{(\mu,\nu)}$. The coefficients $a_{n},b_{n},c_{n},d_{n}$ are the sum of the dimensions of the irreducible representations for the Lie algebra. The modular invariant partition function of $\mathcal{T}^{(\mu,\nu)}$ on torus is
$Z=|\chi_{0}|^{2}+\sum_{i=1}^{3}\mathcal{M}_{i}|\chi_{i}|^{2},$ (40)
where $\mathcal{M}_{i}$ are natural numbers that capture the degeneracy of primaries fields.
The first main result of this paper is as follows. We find that the characters $\chi_{i}$ of the theory $\mathcal{T}^{(\mu,\nu)}$ satisfy the following uniform fourth-order modular linear differential equation
$[D^{4}+\lambda_{1}E_{4}D^{2}+\lambda_{2}E_{6}D+\lambda_{3}E_{4}^{2}]\chi=0,$ (41)
where the three coefficients $\lambda_{1},\lambda_{2}$ and $\lambda_{3}$ are characterized by $(\mu,\nu)$ as follows:
$\lambda_{1}$ $=-\frac{11\mu^{2}\nu^{2}+4\mu^{2}\nu+11\mu^{2}+4\mu\nu^{2}-28\mu\nu+4\mu+11\nu^{2}+4\nu+11}{36(\mu+1)^{2}(\nu+1)^{2}},$ (42)
$\lambda_{2}$ $=\frac{49\mu^{2}\nu^{2}+8\mu^{2}\nu-5\mu^{2}+8\mu\nu^{2}-56\mu\nu+8\mu-5\nu^{2}+8\nu+49}{216(\mu+1)^{2}(\nu+1)^{2}},$ (43)
$\lambda_{3}$ $=-\frac{(\mu\nu-1)^{2}(\mu\nu+2\mu+1)(\mu\nu+2\nu+1)}{16(\mu+1)^{4}(\nu+1)^{4}}.$ (44)
Taking $\nu=5$, the 4th order MLDE factorizes to
$\left(D^{2}-\frac{(\mu-5)(\mu+7)}{144(\mu+1)^{2}}E_{4}\right)\left(D^{2}-\frac{(5\mu-1)(7\mu+1)}{144(\mu+1)^{2}}E_{4}\right)\chi=0.$ (45)
The differential operator can be written as $\Delta_{1/\mu}\circ\Delta_{\mu}$ where $\Delta_{\mu}$ is the second order operator given in (14). This factorization demonstrates that the fourth-order MLDE (41) provides a natural generalization of the second-order Mathur–Mukhi–Sen equation.
It is natural to postulate that the above expressions for the central charge, conformal weights, and the modular linear differential equation (41) continue to hold for the extended magic triangle. This assumption allows us to identify two-dimensional rational conformal field theories beyond the original magic triangle of ordinary Lie groups. Indeed, upon specializing to $\nu=4$, the equation (41) reduces to a one-parameter family of fourth-order MLDEs governing minimal $\mathcal{W}$-algebras Kawasetsu and Sakai (2018). In this case, the characters of the theory $\mathcal{T}^{(\mu,4)}$ coincide with the characters of Ramond-twisted irreducible modules of the minimal $\mathcal{W}$-algebra $\mathcal{W}{-h^{\vee}/6}(\mathfrak{g},f\theta)$, where $\mathfrak{g}$ belongs to the Cvitanovi’c–Deligne exceptional series. More generally, by explicitly solving the MLDE (41) and interpreting its solutions as RCFT characters, we are led to identify a class of theories $\mathcal{T}^{(\mu,\nu)}$ extending those associated with the original magic triangle. [^1] Note that two theories ${\mathcal{T}}^{(\frac{2}{3},2)}$ and ${\mathcal{T}}^{(1,\frac{3}{2})}$ have different central charges $4/5$ and $6/5$ respectively, even though they have the same dimensions 0 and dual Coxeter numbers 1-1.
  • $\mathcal{T}^{(2,1)}$ is a unitary RCFT with $c=2$ and conformal weights $h_{i}=0,(\frac{1}{6})_{6},(\frac{1}{2})_{3},(\frac{2}{3})_{2}$. The VOA was constructed in Kitazume et al. (2002), see Remark 5.8 there for all its irreducible modules. The RCFT is denoted as $M_{2}$ in (Duan et al., 2022, Section 5.6). The four characters have appeared as CFT-type solutions of fourth order holomorphic MLDE in (Kaidi et al., 2021, Table 6). The $q$-series of the characters can be found in (Duan et al., 2022, Equation (5.34)). The $M_{2}$ can be realized by the coset construction $(D_{4})_{1}/(A_{2})_{1}$ (Duan et al., 2022, Section 5.6). We notice that it can also be realized as a non-diagonal modular invariant of $(U_{1})_{2}\times(U_{1})_{6}$, which is consistent with magic triangle element $U_{1}^{2}$.
  • $\mathcal{T}^{(3/2,1)}=D_{2\rm A}$ is a unitary RCFT associated to the 2A conjugacy class of the Monster group. It has $c=6/5$ and conformal weights $h_{i}=0,(\frac{1}{10})_{3},(\frac{1}{2})_{3},\frac{3}{5}$. The VOA can be constructed using a code in $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ and as an extension of the product of Ising model $\mathrm{Vir}_{4,3}$ and tri-critical Ising model $\mathrm{Vir}_{5,4}$ Lam and Yamada (2000). For 2d RCFTs associated to various conjugacy classes and sporadic groups, see a good review in Bae et al. (2021b).
  • $\mathcal{T}^{(3,2/3)}=\mathrm{Vir}_{6,5}^{\rm e}$ is a non-diagonal modular invariant of minimal model $\mathrm{Vir}_{6,5}$ which describes the 3-state Potts model. The VOA can be constructed using a code in $\mathbb{Z}_{3}$ Miyamoto (2001).
  • $\mathcal{T}^{(3/2,3/2)}=(A_{1})_{8}^{\rm e}$ is the D-type modular invariant of $(A_{1})_{8}$. The VOA can be constructed by simple current extension.
All RCFTs associated to the extended magic triangle are presented in Table 2. The data of the extended magic triangle and the RCFTs are collected in the Tables 5, 6, 7 in the Appendix A.
Now we discuss the symmetry of the magic triangle at level one. It is easy to verify that under the exchange of $\mu$ and $\nu$, the central charge $c$, the conformal weight $h_{1}$ and the fourth order MLDE are invariant, while $h_{2},h_{3}$ are exchanged. This suggests that the $\mathcal{T}^{(\mu,\nu)}$ and $\mathcal{T}^{(\nu,\mu)}$ theories are isomorphic to each other, with only the labeling of $\chi_{2}$ and $\chi_{3}$ is exchanged. Therefore we only need to present half of the triangle in the Tables 5, 6, 7. In total, there are 30 non-equivalent RCFTs in the extended magic triangle at level one. The 16 of them come from ordinary WZW model of semisimple Lie groups, 8 of them come from intermediate CFTs, three of them from Virasoro minimal models, two of them from $U(1)$ theory, i.e. compact bosons on a circle, and the last one $D_{2\rm A}$ associated to Monster group. For all 30 theories on the magic triangle, the Fourier coefficients of the characters remain non-negative integers.
A few remarks are in order:
  • We find that the four characters of each theory $\mathcal{T}^{(\mu,\nu)}$ satisfy the following remarkable identity:
    $\chi_{0}\chi_{1}-\chi_{2}\chi_{3}=\mathrm{const}=R_{1}.$ (46)
    We refer to (46) as the magic bilinear relation. Its derivation will be presented in a later section. As an immediate consequence, this identity implies a simple relation among the dimensions $R_{1},R_{2},R_{3}$ of the corresponding representations,
    $\frac{R_{2}R_{3}}{R_{1}}=\frac{12(\mu\nu-1)^{2}}{\mu\nu}.$ (47)
  • For the Cvitanović–Deligne exceptional series, there exist only two distinct characters, or even a single character in the case of $(E_{8})_{1}$. We adopt the convention that the theory $\mathcal{T}^{(5,\nu)}$ has $\chi_{0},\chi_{2}$ as a row, while $\mathcal{T}^{(\mu,5)}$ has $\chi_{0},\chi_{3}$ as a column. In these cases, the remaining solutions of the fourth-order MLDE exhibit Fourier coefficients with unbounded denominators and therefore cannot be interpreted as RCFT characters. Such solutions are referred to as spurious solutions. A detailed discussion of this phenomenon is deferred to Appendix C.
  • The thirty theories $\mathcal{T}^{(\mu,\nu)}$ identified here do not exhaust all CFT-type solutions of general fourth-order holomorphic MLDEs. For instance, the characters of the minimal model $\mathrm{Vir}_{9,2}$ also satisfy a fourth-order holomorphic MLDE. For a systematic classification of admissible solutions with no integer non-vacuum conformal weights, we refer the reader to Kaidi et al. (2021).
3.2 Magic coset
We find that the coset relation (28) within the Cvitanovi’c–Deligne exceptional series is in fact a very special instance of a much broader class of coset constructions that exist among all theories in the extended magic triangle. Remarkably, for any two theories in the triangle one may form a coset that yields a third theory, which again belongs to the same triangle. More precisely, we uncover the following universal coset relation valid throughout the extended magic triangle:
$\frac{\mathcal{T}^{(\mu,\rho)}}{\mathcal{T}^{(\nu,\rho)}}=\mathcal{T}^{(\mu,1/\nu)},\qquad\mu>\nu.$ (48)
We refer to (48) as the magic coset. Using the central charge formula (33), it is straightforward to verify that both sides of (48) indeed have identical central charges. This relation suggests that taking the coset of any upper and lower exceptional series produces another theory within the magic triangle. As concrete examples, consider $(\mu,\nu)=(5,3)$ and $(3,2)$. The magic coset (48) yields
$\frac{(\mathfrak{g}_{\rm CD})_{1}}{(\mathfrak{g}_{\rm sub})_{1}}=(A_{1})_{1},\quad\textrm{and}\quad\frac{(\mathfrak{g}_{\rm sub})_{1}}{(\mathfrak{g}_{\rm Severi})_{1}}=(U_{1})_{3}.$ (49)
These include well-known coset constructions such as $(E_{8})_{1}/(E_{7})_{1}=(A_{1})_{1}$ and $(E_{7})_{1}/(E_{6})_{1}=(U_{1})_{3}$. Similarly, for $(\mu,\nu)=(5,4)$ and $(4,3)$, the magic coset gives
$\frac{(\mathfrak{g}_{\rm CD})_{1}}{(\mathfrak{g}_{\rm I})_{1}}=\mathrm{Vir}^{\rm eff}_{5,2},\quad\textrm{and}\quad\frac{(\mathfrak{g}_{\rm I})_{1}}{(\mathfrak{g}_{\rm sub})_{1}}=\mathrm{Vir}^{\rm eff}_{5,3},$ (50)
both of which were previously observed in Lee et al. (2024b). Equation (48) indicates that by taking cosets among exceptional series one can generate every theory in the extended magic triangle at level one.
An equivalent formulation of the magic coset relation (48) is
$\mathcal{T}^{(\mu,\nu)}=\frac{\mathcal{T}^{(\mu,\rho)}}{\mathcal{T}^{(1/\nu,\rho)}}=\frac{\mathcal{T}^{(\rho,\nu)}}{\mathcal{T}^{(\rho,1/\mu)}}.$ (51)
Setting $\rho=5$ yields
$\mathcal{T}^{(\mu,\nu)}=\frac{\mathcal{T}^{(\mu,5)}}{\mathcal{T}^{(1/\nu,5)}}=\frac{\mathcal{T}^{(5,\nu)}}{\mathcal{T}^{(5,1/\mu)}}.$ (52)
This form is particularly useful, as it implies that every theory in the extended magic triangle can be realized as a coset between two theories belonging to the Cvitanović–Deligne exceptional series. In this sense, the magic coset construction provides a CFT analogue of the centralizer construction of Deligne–Gross for the original magic triangle Deligne and Gross (2002).
Recall the degeneracy of non-vacuum primary of $\mathcal{T}^{(5,\nu)}$ theory is denoted as $\mathcal{N}_{(\nu)}$, which satisfies $\mathcal{N}_{(\nu)}=\mathcal{N}_{(1/\nu)}.$ The coset (52) implies that the degeneracies of $\mathcal{T}^{(\mu,\nu)}$ theory are
$\mathcal{M}^{(\mu,\nu)}_{1,2,3}=\left(\mathcal{N}_{(\mu)}\mathcal{N}_{(\nu)},\mathcal{N}_{(\nu)},\mathcal{N}_{(\mu)}\right).$ (53)
This shows that $\mathcal{M}_{2}$ remains constant along each column, while $\mathcal{M}_{3}$ remains constant along each row of the magic triangle at level one, as summarized in Table 7.
The magic coset (52) further determines the modular $S$-transformation of $\mathcal{T}^{(\mu,\nu)}$. While one might naively expect the solutions of the fourth-order MLDE to form a degree-four vector-valued modular form, it is more natural in the present context to regard the four characters as forming a $2\times 2$ matrix-valued modular form. [^2] We thank Don Zagier for suggesting the interpretation in terms of matrix-valued modular forms. Let us define the $2\times 2$ matrix $X^{(\mu,\nu)}$ for the characters $\chi_{i}$ of $\mathcal{T}^{(\mu,\nu)}$ theory as
$X^{(\mu,\nu)}=\begin{pmatrix}\chi_{0}&\chi_{2}\\ \chi_{3}&\chi_{1}\end{pmatrix}$ (54)
We then find the following elegant $S$-transformation law:
$X^{(\mu,\nu)}(-1/\tau)=S_{(\mu)}\cdot X^{(\mu,\nu)}(\tau)\cdot S_{(\nu)}^{T},$ (55)
where $S_{(\nu)}$ denotes the $2\times 2$ $S$-matrix of the theory $\mathcal{T}^{(5,\nu)}$ in the Cvitanović–Deligne exceptional series. Taking the determinant of (55) and using $\det S_{(\nu)}=-1$, we see that $\det(X^{(\mu,\nu)})=\chi_{0}\chi_{1}-\chi_{2}\chi_{3}$ is modular invariant. Moreover, since $\alpha_{0}+\alpha_{1}=0$ and $\alpha_{2}+\alpha_{3}=1$, this determinant is holomorphic and must therefore be constant. A direct computation shows that the constant equals $R_{1}$, thereby proving the magic bilinear relation (46).
The magic coset construction also leads to a variety of nontrivial relations among characters of level-one theories in the extended magic triangle. In particular, (52) implies the following relations between the characters of the Cvitanović–Deligne exceptional series and those of a general theory $\mathcal{T}^{(\mu,\nu)}$:
$\left(\begin{array}[]{lr}\chi_{0}&\chi_{2}\end{array}\right)^{(5,\nu)}$ $=\left(\begin{array}[]{cc}\chi_{0}&\chi_{2}\end{array}\right)^{(5,1/\mu)}\cdot\left(\begin{array}[]{cc}1&0\\ 0&\mathcal{N}_{(\mu)}\end{array}\right)\cdot X^{(\mu,\nu)},$ (60)
$\left(\begin{array}[]{c}\chi_{0}\\ \chi_{3}\end{array}\right)^{(\mu,5)}$ $=X^{(\mu,\nu)}\cdot\left(\begin{array}[]{cc}1&0\\ 0&\mathcal{N}_{(\nu)}\end{array}\right)\cdot\left(\begin{array}[]{c}\chi_{0}\\ \chi_{3}\end{array}\right)^{(1/\nu,5)}.$ (67)
Extracting the leading terms yields the following relations among representation dimensions:
$R_{1}^{(\mu,\nu)}=\frac{\mu\nu}{12(5\mu-1)(5\nu-1)}R_{(\mu)}R_{(\nu)},$ (68)
$R_{2}^{(\mu,\nu)}=\frac{\mu\nu-1}{5\nu-1}R_{(\nu)},$ (69)
$R_{3}^{(\mu,\nu)}=\frac{\mu\nu-1}{5\mu-1}R_{(\mu)}.$ (70)
Here $R_{(\nu)}$ denotes the dimension of the fundamental representation of the Cvitanović–Deligne exceptional series, cf. (2.3). These relations are readily checked to be consistent with (47). For the row $\mathcal{T}^{(5,\nu)}$, the expression for $R_{3}^{(\mu,\nu)}$ becomes linear in $\nu$ and takes the form
$E_{7\frac{1}{2}}:\ 12\nu-3,\ E_{7}:12\nu-4,\ E_{6}:6\nu-3,\ F_{4}:6\nu-4$ (71)
$D_{4}:2\nu-2,\ G_{2}:2\nu-3,\ A_{2}:\nu-2,\ A_{1}:\nu-3,\ A_{\frac{1}{2}}:\nu-4.$ (72)
This precisely reproduces the extension of the preferred representation dimensions found by Deligne–Gross (Deligne and Gross, 2002, Equation 2).
More generally, the magic coset relation (51) implies the following character identity among three theories in the magic triangle:
$X^{(\mu,\nu)}$ $=X^{(\mu,\rho)}\left(\begin{array}[]{cc}1&0\\ 0&\mathcal{N}_{(\rho)}\end{array}\right)X^{(1/\rho,\nu)},$ (75)
for $\rho\leq\nu$ and $1/\rho\leq\mu$. At leading order, this yields the following relations among representation dimensions:
$d_{\mathfrak{g}}^{(\mu,\nu)}$ $=d_{\mathfrak{g}}^{(\mu,\rho)}+d_{\mathfrak{g}}^{(1/\rho,\nu)}+\mathcal{N}_{(\rho)}R_{2}^{(\mu,\rho)}R_{3}^{(1/\rho,\nu)},$ (76)
$R_{1}^{(\mu,\nu)}$ $=\mathcal{N}_{(\rho)}R_{1}^{(\mu,\rho)}R_{1}^{(1/\rho,\nu)},$ (77)
$R_{2}^{(\mu,\nu)}$ $=R_{2}^{(1/\rho,\nu)}+\mathcal{N}_{(\rho)}R_{2}^{(\mu,\rho)}R_{1}^{(1/\rho,\nu)},$ (78)
$R_{3}^{(\mu,\nu)}$ $=R_{3}^{(\mu,\rho)}+\mathcal{N}_{(\rho)}R_{1}^{(\mu,\rho)}R_{3}^{(1/\rho,\nu)}.$ (79)
3.3 Five atomic theories
An interesting consequence of the magic coset relation (48) is that all thirty CFTs appearing in the extended magic triangle in Table 2 can be generated from only five atomic theories lying below the line $\mu\nu=1$. These five atoms are
$\mathrm{Vir}_{5,2}^{\rm eff},\ \mathrm{Vir}_{5,3}^{\rm eff},\ U(1)_{3},\ \mathrm{Vir}_{6,5}^{\rm e},\ D_{2\rm A}.$ (80)
This means that every theory $\mathcal{T}^{(\mu,\nu)}$ can be realized as an extension of a product of these atomic theories, or equivalently as a non-diagonal modular invariant of an tensor product CFT. More concretely, given a theory $\mathcal{T}^{(\mu,\nu)}$ in Table 2, we draw a right triangle bounded by the three lines $\mu=\mathrm{const}$, $\nu=\mathrm{const}$, and $\mu\nu=1$. The theory $\mathcal{T}^{(\mu,\nu)}$ is then constructed from the atomic theories that lie along the segment of the slope just below the line $\mu\nu=1$ within this triangle. Since each of the five atoms appears twice along this line, any theory in the extended magic triangle involves at most two identical atomic constituents. The basic data of these five atoms are summarized in Table 4.
[…]
$(\mu,\nu)$ $(1/4,5)$ $(1/3,4)$ $(1/2,3)$ $(2/3,2)$ $(1,3/2)$
$\dim$ 1 1 1 0 0
$h^{\vee}$ $3/2$ $2/3$ 0 1-1 1-1
$c$ 2/5 3/5 1 4/5 6/5
$h_{1},R_{1}$ - $1/20,1$ $(1/12)_{2},1$ $(1/15)_{2},1$ $(1/10)_{3},1$
$h_{2},R_{2}$ $1/5,1$ $1/4,1$ $(1/3)_{2},1$ $2/5,1$ $(1/2)_{3},1$
$h_{3},R_{3}$ - $4/5,1$ $3/4,2$ $(2/3)_{2},1$ $3/5,2$
Table 4: The five atoms.
We first illustrate this construction for the Cvitanović–Deligne exceptional series at level one using the coset relation (52). Starting from the bottom-left corner of Table 2, we observe that $\mathcal{T}^{(5,1/4)}={\rm Vir}_{5,2}^{\rm eff}$ is itself one of the atomic theories. The next theory, $\mathcal{T}^{(5,1/3)}=(A_{1})_{1}$ WZW model is obtained by adjoining $\mathcal{T}^{(4,1/3)}={\rm Vir}_{5,3}^{\rm eff}$ to ${\rm Vir}_{5,2}^{\rm eff}$. Similarly, the third theory, $\mathcal{T}^{(5,1/2)}=(A_{2})_{1}$ WZW model is obtained by adjoining $\mathcal{T}^{(3,1/2)}=(U_{1})_{3}$ to $(A_{1})_{1}$. Proceeding in this manner, one successively constructs the level-one WZW models of the Cvitanović–Deligne exceptional series by accumulating diagonal elements. In particular, combining the first five atoms, from $\mathcal{T}^{(5,1/4)}={\rm Vir}_{5,2}^{\rm eff}$ to $\mathcal{T}^{(3/2,1)}=D_{2\rm A}$ yields the $(D_{4})_{1}$ WZW model, while taking two copies of each of these five atoms produces the $(E_{8})_{1}$ WZW model.
In summary, any theory $\mathcal{T}^{(\mu,\nu)}$ listed in Table 2 can be obtained by combining all atomic theories lying along the skew diagonal that runs from the leftmost atom in the $\mu$-th row to the uppermost atom in the $\nu$-th column. For instance, $\mathcal{T}^{(1,2)}=M_{2}$ is constructed from ${\rm Vir}^{\rm e}_{6,5}$ and $D_{2\rm A}$. The $\mathcal{T}^{(3,3/2)}=(C_{3})_{1}$ is built of $(U_{1})_{3},{\rm Vir}_{6,5}^{\rm e}$ and two $D_{2\rm A}$’s.
3.4 The diagonal exceptional series
A distinguished specialization of the magic coset relation (48) is obtained by setting $\rho=\mu$, which yields
$\mathcal{T}^{(\mu,1/\nu)}=\frac{\mathcal{T}^{(\mu,\mu)}}{\mathcal{T}^{(\mu,\nu)}},\qquad\mu>\nu.$ (81)
This relation provides a direct generalization of the $E_{8}$ dual pairs appearing in (28). For instance, when $\mu=3$, equation (81) implies that the RCFTs in the subexceptional series at level one naturally pair via coset constructions with respect to $(D_{6})_{1}$. Given the special interest on $\mathcal{T}^{(\mu,\mu)}$ theories which are
$(A_{1})_{8}^{\rm e},\ (A_{2})_{1}^{2},\ (D_{6})_{1},\ (D_{6+1/2+1/2})_{1},\ (E_{8})_{1},$
we call them the diagonal exceptional series. Each of these theories admits a $\mathbb{Z}_{2}$ inner automorphism such that $h_{2}=h_{3}$. Consequently, the theory $\mathcal{T}^{(\mu,\mu)}$ possesses three independent characters with conformal weights $h_{i}=0,$ $(\mu-1)/(\mu+1)$, $\mu/(\mu+1)$. These characters satisfy a third-order holomorphic modular linear differential equation of the form
$D^{3}+\lambda_{1}E_{4}D+\lambda_{2}E_{6}]\chi=0,$ (82)
with
$\lambda_{1}=-\frac{11\mu^{2}-14\mu+11}{36(\mu+1)^{2}},\quad\lambda_{2}=\frac{(\mu-1)^{2}}{8(\mu+1)^{2}}.$ (83)
The relation (81) gives rise to a variety of nontrivial coset constructions. Here we record two particularly simple examples, which to our knowledge have not appeared previously in the literature:
$D_{2\rm A}$ $=\frac{(A_{1})_{8}^{\rm e}}{D_{2\rm A}},$ (84)
$M_{2}$ $=\frac{(A_{2})_{1}^{2}}{M_{2}}.$ (85)
Using the general character identity (75), one readily derives the corresponding character relations. For the coset (84), one finds
$\chi^{(A_{1})_{8}^{\rm e}}_{0}=\chi^{(A_{1})_{8}}_{0}+\chi^{(A_{1})_{8}}_{2}=\left(\chi^{D_{2\rm A}}_{0}\right)^{2}+3\Big(\chi^{D_{2\rm A}}_{\frac{1}{2}}\Big)^{2}.$ (86)
Here, and in what follows, the subscript of $\chi$ denotes the conformal weight of the corresponding character. Similarly, for the coset (85), the character relation takes the form
$\left(\chi^{(A_{2})_{1}}_{0}\right)^{2}$ $=\left(\chi^{M_{2}}_{0}\right)^{2}+3\Big(\chi^{M_{2}}_{\frac{1}{2}}\Big)^{2}.$ (87)
4 Magic triangle at level two
At level two, the magic triangle no longer behaves as perfect as at level one. For the 18 elements belonging to the original magic triangle in Table 1, the associated higher-level theories present no conceptual difficulties, and the corresponding RCFTs are well defined. By contrast, the remaining 12 theories in the extended triangle require more careful treatment at higher levels. On the one hand, the two elements with vanishing dimension $d=0$ and dual Coxeter number $h^{\vee}=-1$ become trivial at higher levels, since their central charges vanish identically. On the other hand, some intermediate Lie algebras do not admit consistent higher-level generalizations. A typical example is the $E_{7\frac{1}{2}}$ case, for which we have argued that only the levels $k\leq 5$ can be realized as RCFTs Lee et al. (2024a).
Using the dimension and dual Coxeter number formulas of the magic triangle, one obtains a universal expression for the level-two central charge,
$c_{2}=\frac{24(\mu\nu-1)(\mu\nu+\mu+\nu-4)}{(\mu+1)(\nu+1)(\mu\nu+\mu+\nu-3)}$ (88)
In principle, all 13 conformal weights can be written explicitly as functions of $\mu$ and $\nu$. In particular, there is always a vacuum character with exponent zero. We expect that the characters of all level-two theories satisfy a two-parameter family of holomorphic modular linear differential equations of order 13 with index $l=36$. Owing to their complexity, we do not present these equations explicitly. Nevertheless, substantial simplifications occur for certain exceptional series. For example, the characters of the CD exceptional series at level two satisfy a one-parameter family of sixth-order holomorphic MLDEs Lee et al. (2024a). Moreover, the intermediate exceptional series at level two are realized by minimal $\mathcal{W}$-algebras $\mathcal{W}_{-h^{\vee}/6+1}(\mathfrak{g}_{\rm CD},f_{\theta})$, which has been studied recently in Sun (2024); Arakawa et al. (2024); Kac and Wakimoto (2025). Finally, for the subexceptional series at level two, an emergent supersymmetry plays a crucial role. In this case, it is most natural to fermionize the theories and study their fermionic characters and fermionic MLDEs. A detailed analysis of this structure will be presented in the next section. Finally, we expect that the Severi exceptional series at level two should admit a $\mathbb{Z}_{3}$ parafermionic structure. It would be interesting to investigate this in the future.
4.1 Subexceptional series at level two
A remarkable phenomenon we observe is that all theories in subexceptional series $(\mu,\nu)=(3,\nu)$ at level two are supersymmetric. These theories $\mathcal{T}_{2}^{(3,\nu)}$ are
$(U_{1})_{6},\ (A_{1})_{6},\ (A_{1}^{\oplus 3})_{2},\ (C_{3})_{2},\ (A_{5})_{2},\ (D_{6})_{2},\ (D_{6+1/2})_{2},\ (E_{7})_{2}.$ (89)
A prerequisite of emergent supersymmetry is that the RCFT has a primary field with conformal weight $3/2$, which serves as the supersymmetry generator. It is easy to verify this is true for all theories in (88). In many cases, the supersymmetry of these theories has been studied, see e.g. Johnson-Freyd (2021); Bae and Lee (2021). In particular, the $(D_{6+1/2})_{2}$ was found to be supersymmetric recently in Lee et al. (2024b). Taking $\mu=3$ in (88), the central charge of $\mathcal{T}_{2}^{(3,\nu)}$ theory is
$c=\frac{3(3\nu-1)(4\nu-1)}{2\nu(\nu+1)}.$ (90)
We observe that the fermionzation of each theory in (89) denoted as $\mathcal{F}\mathcal{T}_{2}^{(3,\nu)}$ has at most four distinct Neveu–Schwarz characters and equal number of Ramond characters. In general, we find the four Neveu–Schwarz weights are
$0,\frac{4\nu-1}{4(\nu+1)},\frac{2\nu-1}{2\nu},\frac{7\nu^{2}-7\nu+6}{4\nu(\nu+1)},$ (91)
while the four Ramond weights are
$\frac{(3\nu-1)(4\nu-1)}{16\nu(\nu+1)},\frac{3(4\nu-1)}{16\nu},\frac{(3\nu-1)(8\nu+3)}{16\nu(\nu+1)},\frac{3(8\nu-5)}{16\nu}.$ (92)
Notice the first Ramond weight has exponent 0 which indicates the existence of the constant $\widetilde{\rm R}$ character. One can also easily recover 13 bosonic conformal weights and check that the 13 conformal weights always yield a 13th order MLDE with index $l=36$. For theories with less than four distinct NS characters, the fermionic weights take value in the above list.
We further find a one-parameter family of holomorphic fermionic MLDE of degree four for the Neveu–Schwarz characters of $\mathcal{F}\mathcal{T}_{2}^{(3,\nu)}$ theories. To present the fermionic MLDE, we first recall the Jacobi theta functions
$\theta_{2}=\sum_{n\in\mathbb{Z}}q^{(n+\frac{1}{2})^{2}/2},\quad\theta_{3}=\sum_{n\in\mathbb{Z}}q^{n^{2}/2},\quad\theta_{4}=\sum_{n\in\mathbb{Z}}(-1)^{n}q^{n^{2}/2}.$ (93)
The most general form of fermionic MLDE with degree four and index $l=0$ for Neveu–Schwarz characters is Duan et al. (2023)
$\Big[$ $D^{4}+\lambda_{1}(\theta_{4}^{4}-\theta_{2}^{4})D^{3}+\big(\lambda_{2}(\theta_{2}^{8}+\theta_{4}^{8})+\lambda_{3}\theta_{2}^{4}\theta_{4}^{4}\big){D}^{2}+\big(\lambda_{4}(\theta_{4}^{8}\theta_{2}^{4}-\theta_{2}^{8}\theta_{4}^{4})$ (94)
$+\lambda_{5}(\theta_{4}^{12}-\theta_{2}^{12})\big)D+\lambda_{6}(\theta_{2}^{16}+\theta_{4}^{16})+\lambda_{7}(\theta_{4}^{12}\theta_{2}^{4}+\theta_{2}^{12}\theta_{4}^{4})+\lambda_{8}\theta_{2}^{8}\theta_{4}^{8}\Big]\chi^{\rm NS}=0.$
In the meantime, the Ramond characters should satisfy the following dual MLDE
$\Big[D^{4}$ $-\lambda_{1}(\theta_{3}^{4}+\theta_{4}^{4})D^{3}+\big(\lambda_{2}(\theta_{4}^{8}+\theta_{3}^{8})-\lambda_{3}\theta_{4}^{4}\theta_{3}^{4}\big)D^{2}+\big(\lambda_{4}(\theta_{3}^{8}\theta_{4}^{4}+\theta_{4}^{8}\theta_{3}^{4})$ (95)
- $\,\lambda_{5}(\theta_{3}^{12}+\theta_{4}^{12})\big)D+\lambda_{6}(\theta_{4}^{16}+\theta_{3}^{16})-\lambda_{7}(\theta_{3}^{12}\theta_{4}^{4}+\theta_{4}^{12}\theta_{3}^{4})+\lambda_{8}\theta_{4}^{8}\theta_{3}^{8}\Big]\chi^{\rm R}=0.$
The eight fermionic weights (91) and (92) along with the no free fermion condition allow us to solve the unfixed coefficients $\lambda_{i}$ are rational functions of $\nu$. In the end, we find the one-parameter family as
$\lambda_{1}=\frac{\nu^{2}+3\nu-3}{4\nu(\nu+1)},\ \ \lambda_{2}=-\frac{856\nu^{4}-1654\nu^{3}+2827\nu^{2}-2610\nu+1053}{1152\nu^{2}(\nu+1)^{2}},$ (96)
$\lambda_{3}=-\frac{860\nu^{4}-1070\nu^{3}+3047\nu^{2}-3546\nu+1377}{576\nu^{2}(\nu+1)^{2}},$
$\lambda_{4}=\frac{2560\nu^{6}+1488\nu^{5}-13605\nu^{4}+22117\nu^{3}-14274\nu^{2}+4401\nu-675}{9216\nu^{3}(\nu+1)^{3}},$
$\lambda_{5}=\frac{6896\nu^{6}+6816\nu^{5}-33111\nu^{4}+60047\nu^{3}-48078\nu^{2}+17955\nu-2889}{27648\nu^{3}(\nu+1)^{3}},$
$\lambda_{6}=-\frac{(3\nu-1)(4\nu-1)^{2}\left(4\nu^{2}+15\nu-9\right)\left(16\nu^{2}-21\nu+23\right)}{65536\nu^{4}(\nu+1)^{3}},$
$\lambda_{7}=\frac{3(3\nu-1)(4\nu-1)\left(32\nu^{6}-740\nu^{5}+1740\nu^{4}-2133\nu^{3}+1121\nu^{2}-207\nu+27\right)}{16384\nu^{4}(\nu+1)^{4}},$
and $\lambda_{8}=2\lambda_{7}-2\lambda_{6}$. The last equation comes from the supersymmetry condition that one Ramond weight has zero exponent. We checked the Neveu–Schwarz and Ramond characters of $\mathcal{F}\mathcal{T}_{2}^{(3,\nu)}$ theories indeed satisfy the above fermionic MLDEs. This one-parameter family of fermionic MLDE suggests that all Neveu–Schwarz and Ramond characters of $\mathcal{F}\mathcal{T}_{2}^{(3,\nu)}$ theory can be solved as functions of $\nu$ in a uniform way.
4.2 Coset constructions
At level two, there no longer exists a universal coset construction covering the entire magic triangle, in contrast to the level-one case. Nevertheless, the magic coset relation (48) continues to provide strong guidance and motivates us to search for new uniform coset constructions involving the Cvitanović–Deligne, sub and Severi exceptional series at level two.
First we find the following universal coset construction between Cvitanović–Deligne and sub exceptional series at level two:
$\frac{\mathcal{T}^{(5,\nu)}_{2}}{\mathcal{T}^{(3,\nu)}_{2}\times\mathcal{T}^{(5,1/3)}_{2}}=\frac{(\mathfrak{g}_{\rm CD})_{2}}{(\mathfrak{g}_{\rm sub})_{2}\times(A_{1})_{2}}=\mathrm{Vir}_{n+1,n},\qquad n=\left\{\begin{array}[]{ll}3+\frac{4}{\nu-1},&\nu>1\\ \frac{4}{1-\nu}-4,&\nu<1.\end{array}\right$. (99)
This coset holds directly for $\nu=\frac{1}{2},\frac{2}{3},\frac{3}{2},2,3,5$. For $\nu=\frac{1}{4},4$, the relation continues to hold after rescaling $n$ and $n+1$ so that they are coprime integers, in which case both sides of the coset become non-unitary. To the best of our knowledge, the uniform coset construction (99), as well as the one in (104) below, are new. As an illustrative example, for $\nu=\frac{2}{3}$, the coset (99) yields
$\frac{(G_{2})_{2}}{(A_{1})_{6}\times(A_{1})_{2}}=\mathrm{Vir}_{9,8}.$ (100)
Each of the four characters of $(G_{2})_{2}$ decomposes to the summation of 12 terms of $(A_{1})_{6}\times(A_{1})_{2}\times\mathrm{Vir}_{9,8}$. For instance, denote $\mathrm{Vir}_{9,8}$ as $M$, then for the vacuum we have
$\chi^{(G_{2})_{2}}_{0}=$ $\ \chi^{(A_{1})_{2}}_{0}\chi^{(A_{1})_{6}}_{0}\chi^{M}_{0}+\chi^{(A_{1})_{2}}_{\frac{3}{16}}\chi^{(A_{1})_{6}}_{\frac{15}{32}}\chi^{M}_{\frac{11}{32}}+\chi^{(A_{1})_{2}}_{0}\chi^{(A_{1})_{6}}_{\frac{3}{4}}\chi^{M}_{\frac{5}{4}}$ (101)
$+\chi^{(A_{1})_{2}}_{\frac{1}{2}}\chi^{(A_{1})_{6}}_{\frac{1}{4}}\chi^{M}_{\frac{5}{4}}+\chi^{(A_{1})_{2}}_{\frac{1}{2}}\chi^{(A_{1})_{6}}_{\frac{3}{2}}\chi^{M}_{0}+\chi^{(A_{1})_{2}}_{0}\chi^{(A_{1})_{6}}_{\frac{3}{4}}\chi^{M}_{\frac{5}{4}}$
$+\chi^{(A_{1})_{2}}_{\frac{3}{16}}\chi^{(A_{1})_{6}}_{\frac{3}{32}}\chi^{M}_{\frac{87}{32}}+\chi^{(A_{1})_{2}}_{\frac{3}{16}}\chi^{(A_{1})_{6}}_{\frac{35}{32}}\chi^{M}_{\frac{87}{32}}+\chi^{(A_{1})_{2}}_{0}\chi^{(A_{1})_{6}}_{\frac{1}{4}}\chi^{M}_{\frac{19}{4}}$
$+\chi^{(A_{1})_{2}}_{\frac{1}{2}}\chi^{(A_{1})_{6}}_{\frac{3}{4}}\chi^{M}_{\frac{19}{4}}+\chi^{(A_{1})_{2}}_{\frac{3}{16}}\chi^{(A_{1})_{6}}_{\frac{15}{32}}\chi^{M}_{\frac{235}{32}}+\chi^{(A_{1})_{2}}_{0}\chi^{(A_{1})_{6}}_{\frac{3}{2}}\chi^{M}_{\frac{21}{2}}.$
The twelve terms are ordered according to increasing total conformal weight.
We further find the following universal coset construction between the sub and Severi exceptional series at level 2:
$\frac{\mathcal{T}^{(3,\nu)}_{2}}{\mathcal{T}^{(2,\nu)}_{2}\times\mathcal{T}^{(3,1/2)}_{2}}=\frac{(\mathfrak{g}_{\rm sub})_{2}}{(\mathfrak{g}_{\rm Severi})_{2}\times(U_{1})_{6}}=\mathcal{B}\mathrm{Vir}_{n+2,n}^{N=1},\qquad n=\left\{\begin{array}[]{ll}4+\frac{4}{\nu-1},&\nu>1,\\ \frac{4}{1-\nu}-6,&\nu<1.\end{array}\right$. (104)
Here again $\nu=\frac{1}{2},\frac{2}{3},\frac{3}{2},2,3,5$. The $\mathcal{B}\mathrm{Vir}_{n+2,n}^{N=1}$ is the bosonization $\mathcal{B}$ of the unitary $N=1$ supersymmetric minimal model $\mathrm{Vir}_{n+2,n}^{N=1}$. For example, taking $\nu=3$ gives
$\frac{(E_{7})_{2}}{(E_{6})_{2}\times(U_{1})_{6}}=\mathcal{B}\mathrm{Vir}_{7,5}^{N=1}$ (105)
It is easy to find $\mathcal{B}\mathrm{Vir}_{7,5}^{N=1}$ has in total $18$ bosonic primary fields with conformal weights
$\left\{0,\frac{3}{70},\frac{29}{560},\frac{9}{112},\frac{4}{35},\frac{3}{14},\frac{27}{80},\frac{269}{560},\frac{19}{35},\frac{43}{70},\frac{73}{112},\frac{5}{7},\frac{9}{10},\frac{8}{7},\frac{7}{5},\frac{3}{2},\frac{23}{14},\frac{31}{16}\right\}.$
Denote $\mathcal{B}\mathrm{Vir}_{7,5}^{N=1}$ as $\mathcal{B}$, then for the vacuum, we find the character character relation
$\chi^{(E_{7})_{2}}_{0}=$ $\ \chi^{(E_{6})_{2}}_{0}\chi^{(U_{1})_{6}}_{0}\chi^{\mathcal{B}}_{0}+2\chi^{(E_{6})_{2}}_{\frac{13}{21}}\chi^{(U_{1})_{6}}_{\frac{1}{6}}\chi^{\mathcal{B}}_{\frac{3}{14}}+\chi^{(E_{6})_{2}}_{\frac{6}{7}}\chi^{(U_{1})_{6}}_{0}\chi^{\mathcal{B}}_{\frac{8}{7}}$ (106)
$+\chi^{(E_{6})_{2}}_{\frac{9}{7}}\chi^{(U_{1})_{6}}_{0}\chi^{\mathcal{B}}_{\frac{5}{7}}+2\chi^{(E_{6})_{2}}_{\frac{13}{21}}\chi^{(U_{1})_{6}}_{\frac{2}{3}}\chi^{\mathcal{B}}_{\frac{5}{7}}+2\chi^{(E_{6})_{2}}_{\frac{4}{3}}\chi^{(U_{1})_{6}}_{\frac{2}{3}}\chi^{\mathcal{B}}_{0}$
$+2\chi^{(E_{6})_{2}}_{\frac{25}{21}}\chi^{(U_{1})_{6}}_{\frac{1}{6}}\chi^{\mathcal{B}}_{\frac{23}{14}}+2\chi^{(E_{6})_{2}}_{\frac{4}{3}}\chi^{(U_{1})_{6}}_{\frac{1}{6}}\chi^{\mathcal{B}}_{\frac{3}{2}}+2\chi^{(E_{6})_{2}}_{\frac{25}{21}}\chi^{(U_{1})_{6}}_{\frac{2}{3}}\chi^{\mathcal{B}}_{\frac{8}{7}}$
$+\chi^{(E_{6})_{2}}_{0}\chi^{(U_{1})_{6}}_{\frac{3}{2}}\chi^{\mathcal{B}}_{\frac{3}{2}}+\chi^{(E_{6})_{2}}_{\frac{9}{7}}\chi^{(U_{1})_{6}}_{\frac{3}{2}}\chi^{\mathcal{B}}_{\frac{3}{14}}+\chi^{(E_{6})_{2}}_{\frac{6}{7}}\chi^{(U_{1})_{6}}_{\frac{3}{2}}\chi^{\mathcal{B}}_{\frac{23}{14}}.$
We check the degeneracy matches and it is straightforward to upgrade the character relation to module decomposition.
Furthermore, there exists the following universal coset Cheng et al. (2020) for Cvitanović–Deligne exceptional series resembling the Maverick coset
$\frac{\mathcal{T}^{(5,\nu)}_{1}\times\mathcal{T}^{(5,\nu)}_{1}}{\mathcal{T}^{(5,\nu)}_{2}}=\frac{(\mathfrak{g}_{\rm CD})_{1}\times(\mathfrak{g}_{\rm CD})_{1}}{(\mathfrak{g}_{\rm CD})_{2}}=\mathrm{Vir}_{n+1,n},\qquad n=\left\{\begin{array}[]{ll}2+\frac{4}{\nu-1},&\nu>1\\ \frac{4}{1-\nu}-3,&\nu<1.\end{array}\right$. (109)
Here $\nu$ takes value in $1/3,1/2,2/3,3/2,2,3,5$. For $\nu=1/4,4$, it still holds if we scale $n$ and $n+1$ such that they are coprime integers and take the proper non-diagonal modular invariant. For $\nu=4$, the non-diagonal modular invariant was identified in Lee et al. (2024a). For $\nu=1/4$, the coset was recently proved in Feng and Jiang (2025). We newly find a similar coset construction for the subexceptional series:
$\frac{(\mathfrak{g}_{\rm sub})_{1}\times(\mathfrak{g}_{\rm sub})_{1}}{(\mathfrak{g}_{\rm sub})_{2}}=\mathcal{B}\mathrm{Vir}_{n+2,n}^{N=1},\qquad n=\left\{\begin{array}[]{ll}2+\frac{4}{\nu-1},&\nu>1,\\ \frac{4}{1-\nu}-4,&\nu<1.\end{array}\right$. (112)
We have a final remark for the intermediate exceptional series at level 2. For $\nu=2,3,5$ and $n=3+\frac{4}{\nu-1}$, the following cosets were found in (Sun, 2024, Equation (1.9),(1.10)):
$\frac{\mathcal{T}^{(5,\nu)}_{2}}{\mathcal{T}^{(4,\nu)}_{2}}$ $=\frac{(\mathfrak{g}_{\rm CD})_{2}}{(\mathfrak{g}_{I})_{2}}=(\mathrm{Vir}_{3n+1,n})_{\rm eff},$ (113)
$\frac{\mathcal{T}^{(4,\nu)}_{2}}{\mathcal{T}^{(3,\nu)}_{2}}$ $=\frac{(\mathfrak{g}_{I})_{2}}{(\mathfrak{g}_{\rm sub})_{2}}=(\mathcal{B}\mathrm{Vir}^{N=1}_{3n+1,n+1})_{\rm eff}.$ (114)
5 Magic triangle at arbitrary positive integer levels
Denote the RCFT associated with the triangle element $(\mu,\nu)$ at level $k$ as $\mathcal{T}^{(\mu,\nu)}_{k}$. For the 16 elements of the magic triangle that are semisimple Lie algebras $\mathfrak{g}$, the corresponding RCFT with arbitrary positive integer level $k$ are just their WZW models $(\mathfrak{g})_{k}$, except for ${\mathfrak{g}}^{(3,2/3)}=A_{1}$, ${\mathfrak{g}}^{(3/2,3/2)}=A_{1}$ and ${\mathfrak{g}}^{(2,3/2)}=A_{2}$ for which the levels must be rescaled by $3$, 88, and $2$, respectively. For the remaining 14 elements, a case-by-case analysis is required. By carefully examining the central charges and conformal weights, we arrive at the following conclusions:
  • $\mathcal{T}^{(1/4,5)}_{k}$ is the effective minimal model $\mathrm{Vir}_{2k+3,2}^{\rm eff}$, see e.g. Duan et al. (2022). This is denoted as level-$k$ Lee-Yang model $(LY)_{k}$ in Duan et al. (2022).
  • $\mathcal{T}^{(1/3,4)}_{k}$ is the effective minimal model $\mathrm{Vir}_{3k+2,3}^{\rm eff}$. This has been realized in Lee et al. (2024b).
  • $\mathcal{T}^{(1/2,3)}_{k}$ is the compact boson $U(1)_{3k}$ with $c=1$.
  • $\mathcal{T}^{(2/3,2)}_{k}$ and $\mathcal{T}^{(1,3/2)}_{k}$ are trivial for $k>1$ due to $d=0$ and $h^{\vee}=-1$.
  • $\mathcal{T}^{(1,2)}_{k}$ is a non-diagonal modular invariant of $U(1)_{2k}\times U(1)_{6k}$. The vacuum character is given by
    $\chi^{U(1)_{2k}}_{0}\chi^{U(1)_{6k}}_{0}+\chi^{U(1)_{2k}}_{\frac{k}{2}}\chi^{U(1)_{6k}}_{\frac{3k}{2}}.$ (115)
  • $\mathcal{T}^{(3/2,3/2)}_{k}$ is the $D$-type non-diagonal modular invariant of $(A_{1})_{8k}$.
  • $\mathcal{T}^{(4,5)}_{k}$, $\mathcal{T}^{(4,3)}_{k}$, $\mathcal{T}^{(4,2)}_{k}$ exist up to $k=5,3,2$ respectively. These theories are equivalent to the minimal $\mathcal{W}$-algebras of E-type with negative integer levels, which have recently been studied in Sun (2024); Arakawa et al. (2024); Kac and Wakimoto (2025). Other $\mathcal{T}^{(4,\nu)}_{k}$ theories for $k>1$ should not exist.
6 Concluding remarks
In this work, we have extended the magic triangle and identified the two-dimensional conformal field theories associated with its elements at arbitrary positive integer levels. In particular, for level-one theories we uncover a number of remarkable and highly structured properties, including modular linear differential equations, the magic coset construction, five atomic theories, magic bilinear relations, and matrix-valued modularity. For level-two theories, we focus on the subexceptional series and found they are all supersymmetric and their fermionic characters satisfy a uniform fermionic MLDE. We also find many interesting coset constructions involving the Cvitanović–Deligne, sub and Severi exceptional series at level two.
Our analysis in this paper is primarily focused on the characters of these theories. It remains unclear, however, whether one can fully define the conformal field theory with the central charge given in magic triangle, or the correlation functions, for some of the theories. For example it would be interesting to find out whether the two atomic models $\mathrm{Vir}^{\rm eff}_{5,2}$ and $\mathrm{Vir}^{\rm eff}_{5,3}$ are sensible conformal field theories with central charge $2/5$ and $3/5$ with proper Hilbert space and correlation functions.
We have not explored much of the dimension of irreducible representations for the magic triangle in this work. The magic coset we have uncovered suggest that the representations of all algebras appearing in the magic triangle are intricately entangled. It would be interesting to find some simple universal formulas for the dimension and dual Coxeter number of these representations, besides the adjoint representation.
Finally, although our discussion has focused on WZW models at positive integer levels, theories at fractional and negative levels are also of significant interest. For instance, it is well known that the Cvitanović–Deligne exceptional series at level $h^{\vee}/6-1$ gives rise to a celebrated family of quasi-lisse vertex operator algebras, whose vacuum characters satisfy a one-parameter family of second-order MLDEs Beem and Rastelli (2018); Arakawa and Kawasetsu (2018). It is intriguing to consider whether this phenomenon can be generalized to other exceptional series, or even to the entire magic triangle.
Acknowledgements
We thank Chongying Dong, Heeyeon Kim, Ching Hung Lam, Jaewon Song, Haowu Wang, Xin Wang, Yinan Wang and Don Zagier for discussion. KL is supported in part by the BIMSA start-up fund and the Beijing Natural Science Foundation International Scientist Project No.IS25024. KS is supported by NSF No.1250010384.
Appendix A Explicit data of the extended magic triangle
We collect the dual Coxeter number of dimension of each element of the extended magic triangle in Table 5. We collect the dimensions of $R_{1},R_{2},R_{3}$ for all theories at level one in Table 6. We collect the central charge, nonzero conformal weights and degeneracies for all theories at level one in Table 7. Due to the symmetry between $\mu$ and $\nu$, we only show the $\mu\leq\nu$ part of the triangle.
$\{-1,0\}$ $\left\{\frac{1}{4},3\right\}$ $3/2$
$\{-1,0\}$ $\{0,2\}$ $\{\frac{3}{2},8\}$ $\{3,16\}$ 2
$\{0,1\}$ $\left\{\frac{2}{3},3\right\}$ $\{2,9\}$ $\{4,21\}$ $\{6,35\}$ $\{10,66\}$ 3
$\left\{\frac{2}{3},1\right\}$ $\left\{\frac{3}{2},4\right\}$ $\left\{\frac{7}{3},8\right\}$ $\{4,18\}$ $\left\{\frac{13}{2},36\right\}$ $\{9,56\}$ $\{14,99\}$ $\{19,144\}$ 4
$\left\{\frac{3}{2},1\right\}$ $\{2,3\}$ $\{3,8\}$ $\{4,14\}$ $\{6,28\}$ $\{9,52\}$ $\{12,78\}$ $\{18,133\}$ $\{24,190\}$ $\{30,248\}$ 5
$1/4$ $1/3$ $1/2$ $2/3$ 11 $3/2$ 22 33 44 55 $\nu\backslash\mu$
Table 5: $\{h^{\vee},\dim\}$ associated to the extended magic triangle.
$\left\{1,1,2\right\}$ $\left\{3,5,5\right\}$ $3/2$
$\left\{1,1,1\right\}$ $\left\{1,2,3\right\}$ $\left\{3,8,6\right\}$ $\left\{3,9,9\right\}$ 22
$\left\{1,1,2\right\}$ $\left\{2,3,4\right\}$ $\left\{2,4,8\right\}$ $\left\{6,14,14\right\}$ $\left\{6,15,20\right\}$ $\left\{12,32,32\right\}$ 33
$\left\{1,1,1\right\}$ $\left\{1,2,3\right\}$ $\left\{2,5,5\right\}$ $\left\{2,6,9\right\}$ $\left\{6,20,15\right\}$ $\left\{6,21,21\right\}$ $\left\{12,44,33\right\}$ $\left\{12,45,45\right\}$ 44
$\left\{1\right\}$ $\left\{2\right\}$ $\left\{3\right\}$ $\left\{7\right\}$ $\left\{8\right\}$ $\left\{26\right\}$ $\left\{27\right\}$ $\left\{56\right\}$ $\left\{57\right\}$ - 55
$1/4$ $1/3$ $1/2$ $2/3$ 11 $3/2$ 2 3 4 5 $\nu\backslash\mu$
Table 6: Dimension $\{R_{1},R_{2},R_{3}\}$ associated to the extended magic triangle at level 11.
$\left\{\frac{6}{5},(\frac{1}{10})_{3},(\frac{1}{2})_{3},\frac{3}{5}\right\}$ $\left\{\frac{12}{5},\frac{1}{5},\frac{3}{5},\frac{3}{5}\right\}$ $3/2$
$\left\{\frac{4}{5},(\frac{1}{15})_{2},\frac{2}{5},(\frac{2}{3})_{2}\right\}\!\!$ $\left\{2,(\frac{1}{6})_{6},(\frac{1}{2})_{3},(\frac{2}{3})_{2}\right\}$ $\left\{\frac{16}{5},(\frac{4}{15})_{2},\frac{3}{5},(\frac{2}{3})_{2}\right\}\!$ $\left\{4,(\frac{1}{3})_{4},(\frac{2}{3})_{2},(\frac{2}{3})_{2}\right\}\!$ 22
$\left\{1,(\frac{1}{12})_{2},(\frac{1}{3})_{2},\frac{3}{4}\right\}$ $\left\{\frac{9}{5},\frac{3}{20},\frac{2}{5},\frac{3}{4}\right\}$ $\left\{3,(\frac{1}{4})_{3},(\frac{1}{2})_{3},\frac{3}{4}\right\}$ $\left\{\frac{21}{5},\frac{7}{20},\frac{3}{5},\frac{3}{4}\right\}$ $\left\{5,(\frac{5}{12})_{2},(\frac{2}{3})_{2},\frac{3}{4}\right\}$ $\left\{6,\frac{1}{2},\frac{3}{4},\frac{3}{4}\right\}$ 33
$\left\{\frac{3}{5},\frac{1}{20},\frac{1}{4},\frac{4}{5}\right\}$ $\left\{\frac{8}{5},(\frac{2}{15})_{2},(\frac{1}{3})_{2},\frac{4}{5}\right\}$ $\left\{\frac{12}{5},\frac{1}{5},\frac{2}{5},\frac{4}{5}\right\}$ $\left\{\frac{18}{5},(\frac{3}{10})_{3},(\frac{1}{2})_{3},\frac{4}{5}\right\}$ $\left\{\frac{24}{5},\frac{2}{5},\frac{3}{5},\frac{4}{5}\right\}$ $\left\{\frac{28}{5},(\frac{7}{15})_{2},(\frac{2}{3})_{2},\frac{4}{5}\right\}$ $\left\{\frac{33}{5},\frac{11}{20},\frac{3}{4},\frac{4}{5}\right\}$ $\left\{\frac{36}{5},\frac{3}{5},\frac{4}{5},\frac{4}{5}\right\}$ 44
$\left\{\frac{2}{5},\underline{\frac{1}{30}},\frac{1}{5},\underline{\frac{5}{6}}\right\}\!\!$ $\left\{1,\underline{\frac{1}{12}},\frac{1}{4},\underline{\frac{5}{6}}\right\}$ $\left\{2,\underline{\frac{1}{6}},(\frac{1}{3})_{2},\underline{\frac{5}{6}}\right\}$ $\left\{\frac{14}{5},\underline{\frac{7}{30}},\frac{2}{5},\underline{\frac{5}{6}}\right\}$ $\left\{4,\underline{\frac{1}{3}},(\frac{1}{2})_{3},\underline{\frac{5}{6}}\right\}$ $\left\{\frac{26}{5},\underline{\frac{13}{30}},\frac{3}{5},\underline{\frac{5}{6}}\right\}$ $\left\{6,\underline{\frac{1}{2}},(\frac{2}{3})_{2},\underline{\frac{5}{6}}\right\}$ $\left\{7,\underline{\frac{7}{12}},\frac{3}{4},\underline{\frac{5}{6}}\right\}$ $\left\{\frac{38}{5},\underline{\frac{19}{30}},\frac{4}{5},\underline{\frac{5}{6}}\right\}\!$ $\left\{8,\underline{\frac{2}{3},\frac{5}{6},\frac{5}{6}}\right\}\!$ 55
$1/4$ $1/3$ $1/2$ $2/3$ 11 $3/2$ 2 3 4 5 $\nu\backslash\mu$
Table 7: CFT data $\{c,h_{1},h_{2},h_{3}\}$ associated to the extended magic triangle at level 11. Those with underlines indicate spurious solutions of the 4th order MLDE.
Appendix B S-matrices
In this section, we collect the $S$-matrices for the two characters $\chi_{0},\chi_{1}$ of the Mathur–Mukhi–Sen series, i.e. the extended Cvitanović–Deligne exceptional series at level one, reviewed in Section 2.3. The $2\times 2$ $S$-matrix $S_{(\nu)}$ for the two characters of $\mathcal{T}^{(\nu,5)}$ which leaves the partition function (15) invariant satisfies $S_{(\nu)}=S_{(1/\nu)}.$ The $S$-matrices for $\nu=1/4,1/3,1/2,2/3,1$ of ${\rm Vir}_{5,2}^{\rm eff},(A_{1})_{1},(A_{2})_{1},(G_{2})_{1},(D_{4})_{1}$ theories are, respectively,
$S_{(4)}=\frac{\varphi^{1/2}}{5^{1/4}}\left(\begin{array}[]{cc}1&\varphi^{-1}\\ \varphi^{-1}&-1\end{array}\right),\quad S_{(3)}=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cr}1&1\\ 1&-1\end{array}\right),\quad S_{(2)}=\frac{1}{\sqrt{3}}\left(\begin{array}[]{cr}1&2\\ 1&-1\end{array}\right),$ (122)
$S_{(3/2)}=\frac{1}{5^{1/4}\varphi^{1/2}}\left(\begin{array}[]{cc}1&\varphi\\ \varphi&-1\end{array}\right),\quad S_{(1)}=\frac{1}{2}\left(\begin{array}[]{cr}1&3\\ 1&-1\end{array}\right),$ (127)
where $\varphi=(\sqrt{5}+1)/2$ is the Golden ratio. The partition function $Z=|\chi_{0}|^{2}+\mathcal{N}|\chi_{1}|^{2}$ is invariant under these $S$ matrices. Note that
$S_{(\nu)}^{T}\cdot\left(\begin{array}[]{cc}1&0\\ 0&\mathcal{N}_{(\nu)}\end{array}\right)\cdot S_{(\nu)}=\left(\begin{array}[]{cc}1&0\\ 0&\mathcal{N}_{(\nu)}.\end{array}\right)$ (132)
In addition for arbitrary $\nu$ we have $\det S_{(\nu)}=-1.$
For the cases with degeneracy, i.e. $\nu=1/2,1,2$, we can have extended $S$-matrices for $\nu=1/2,1$ as
$\frac{1}{\sqrt{3}}\left(\begin{array}[]{ccc}1&1&1\\ 1&w&w^{2}\\ 1&w^{2}&w\end{array}\right),\quad\frac{1}{2}\left(\begin{array}[]{crrr}1&1&1&1\\ 1&1&-1&-1\\ 1&-1&1&-1\\ 1&-1&-1&1\end{array}\right),$ (140)
where $w=e^{2\pi i/3}$. Here all extended $S$-matrices are symmetric and unitary. From (30), the invariance of the pairs leads to the extended $S$-matrices of $\mathcal{T}^{(\nu,5)}$ and $\mathcal{T}^{(1/\nu,5)}$ are related by $S_{(\nu)}=S_{(1/\nu)}^{\dagger}.$ The $S$-matrix of $\mathcal{T}^{(\mu,\nu)}$ in full $(1+\mathcal{N}_{(\mu)})\times(1+\mathcal{N}_{(\nu)})$ form is then
$\left(\begin{array}[]{cc}\chi_{0}&\chi_{2}\\ \chi_{3}&\chi_{1}\end{array}\right)^{(\mu,\nu)}\!\!\!\!(-1/\tau)\ =\ S_{(\mu)}\cdot\left(\begin{array}[]{cc}\chi_{0}&\chi_{2}\\ \chi_{3}&\chi_{1}\end{array}\right)^{(\mu,\nu)}\!\!\!\!(\tau)\cdot S_{(\nu)}^{\dagger}.$ (145)
Appendix C Spurious solutions of MLDE
For each $\mu$, besides the two characters of $\mathcal{T}^{(5,\mu)}$, there are two spurious solutions of the fourth order MLDE (41) with Fourier coefficients with unbounded denominators. For example, for $\mu=1/4$, the two genuine solutions are the Lee–Yang characters $\chi_{0},\chi_{1/5}$, while the two spurious solutions are
$f_{\frac{1}{30}}=$ $\ q^{\frac{1}{60}}\Big(1-\frac{2q}{5}+\frac{q^{2}}{11}+\frac{26q^{3}}{85}+\frac{434q^{4}}{1265}+\frac{9824q^{5}}{27115}+\dots\Big),$
$f_{\frac{5}{6}}=$ $\ q^{\frac{49}{60}}\Big(1+\frac{38q}{33}+\frac{371q^{2}}{561}+\frac{22558q^{3}}{12903}+\frac{383219q^{4}}{374187}+\dots\Big).$
Although these two solutions are not characters themselves, we observe that they are closely related to the characters of the $(E_{7+1/2})_{1}$ theory:
$-180\Delta_{\frac{1}{4}}f_{\frac{1}{30}}=$ $\,\eta^{8}\chi_{0}^{(E_{7+1/2})_{1}},$
$108\Delta_{\frac{1}{4}}f_{\frac{5}{6}}=$ $\,\eta^{8}\chi_{\frac{4}{5}}^{(E_{7+1/2})_{1}}.$
These relations are easy to prove since $\Delta_{4}$ annihilates both sides of the equations. If we define $\chi_{\frac{1}{30}}=R_{1}f_{\frac{1}{30}}$ and $\chi_{\frac{5}{6}}=R_{3}f_{\frac{5}{6}}$, by properly choosing the prefactors such that $R_{3}/R_{1}=3/5$, we find a bilinear relation
$\chi_{0}\chi_{\frac{1}{30}}-\chi_{\frac{1}{5}}\chi_{\frac{5}{6}}=R_{1}.$ (146)
similar to the magic bilinear relation (46). This choice also makes the relation (47) hold. For other $\mu\neq 5$, we obtain similar relations.
The level one $E_{8}$ theory has only one character $\chi_{0}=j^{1/3}(\tau)$ where $j(\tau)$ is the Klein $j$-function. The vacuum character satisfies the first order MLDE,
$\Big(D+\frac{E_{6}}{3E_{4}}\Big)j^{1/3}=0.$ (147)
The second order MLDE (10) for $E_{8}$ theory gets factorized as
$D^{2}-\frac{E_{4}}{6}=\Big(D-\frac{E_{6}}{3E_{4}}\Big)\Big(D+\frac{E_{6}}{3E_{4}}\Big).$ (148)
For $\mu,\nu=5$, the 4th order MLDE (41) becomes
$\left[D^{4}-\frac{1}{6}E_{4}D^{2}+\frac{1}{9}E_{6}D-\frac{1}{36}E_{4}^{2}\right]\chi=0.$ (149)
It factorizes to $D^{2}(D^{2}-\frac{1}{6}E_{4})]\chi=0$. Beside $\chi_{0}=j^{1/3}$ the vacuum character of $(E_{8})_{1}$, this MLDE has two more spurious solutions
$f_{\frac{2}{3}}=$ $\ q^{\frac{1}{3}}\Big(1+\frac{712q}{25}+\frac{75436q^{2}}{275}+\frac{95654016q^{3}}{51425}+\dots\Big),$
$f_{\frac{5}{6}}=$ $\ q^{\frac{1}{2}}\Big(1+\frac{228q}{11}+\frac{34938q^{2}}{187}+\frac{5163352q^{3}}{4301}+\dots\Big).$
These are not characters, but satisfy the following simple relations
$-9\Big(D^{2}-\frac{1}{6}E_{4}\Big)f_{\frac{2}{3}}=$ $\ \eta^{8},$
$\frac{6}{5}\Big(D+\frac{E_{6}}{3E_{4}}\Big)f_{\frac{5}{6}}=$ $\ \eta^{4}j^{-1/3}.$
We find a bilinear relation
$\chi_{0}f_{\frac{2}{3}}-\frac{2^{8}\cdot 3^{3}}{5^{2}}f_{\frac{5}{6}}^{2}=1,$ (150)
similar to the magic bilinear relation (46).

### v2

@@
 The magic triangle due to Cvitanović and Deligne–Gross is an extension of the Freudenthal–Tits magic square of semisimple Lie algebras.
-In this paper, we identify all 2d rational conformal field theories associated to the magic triangle.
+In this paper, we identify all two-dimensional rational conformal field theories associated to the magic triangle.
 These include various Wess–Zumino–Witten (WZW) models, Virasoro minimal models, compact bosons and their non-diagonal modular invariants.
-At level one, we find a two-parameter family of modular linear differential equation of fourth order whose solutions produce the affine characters of all elements in the magic triangle.
-We find a universal coset relation for the whole triangle which generalizes the dual pairs with respect to $(E_{8})_{1}$ in the Cvitanović–Deligne exceptional series.
-This leads to the dimension and degeneracy of each primary field and also to five atomic models which constitute all theories in the triangle.
-At level two, we find a special row of the triangle – the subexceptional series has novel $N=1$ supersymmetry, and the Neveu–Schwarz/Ramond characters satisfy a one-parameter family of fermionic modular linear differential equations.
-Moreover, we find many new coset constructions involving WZW models at higher levels.
+At level one, we uncover a two-parameter family of fourth-order modular linear differential equation whose solutions yield the affine characters of all elements in the magic triangle.
+We further establish a universal coset relation for the whole triangle, generalizing the dual-pair structure with respect to $(E_{8})_{1}$ in the Cvitanović–Deligne exceptional series.
+This coset structure determines the dimensions and degeneracies of all primary fields and leads to five atomic models from which all theories in the triangle can be constructed.
+At level two, we find that a distinghuished row of the triangle – the subexceptional series– exhibits emergent $N=1$ supersymmetry.
+The corresponding Neveu–Schwarz/Ramond characters satisfy a one-parameter family of fermionic modular linear differential equations.
+In addition, we find several new uniform coset constructions involving WZW models at higher levels.
 1 Introduction
@@
 From another perspective, the magic triangle can be regarded as a two-dimensional generalization of the celebrated Cvitanović–Deligne exceptional series Cvitanovic (1981, 2008); Deligne (1996).
-The original magic triangle of Lie groups introduced by Cvitanović (Cvitanovic, 1981, Table 6) is reproduced in Table 1.
-The magic triangle of Deligne and Gross appears in (Deligne and Gross, 2002, Table 1).
-The magic triangle is symmetric, and each row (or equivalently, each column) is referred to as an exceptional series.
+The original magic triangle of Lie groups, found by Cvitanović (Cvitanovic, 1981, Table 6) from certain Diophantine conditions on representation dimensions in the context of extended supergravity theories, is reproduced in Table 1.
+The magic triangle of Deligne and Gross is built as a system of centralizers from dual pairs in $E_{8}$, see (Deligne and Gross, 2002, Table 1).
+The magic triangle is symmetric, and each row (or, equivalently, each column) is referred to as an exceptional series.
 In Table 1, the bottom row, terminating at $E_{8}$, is known as the Cvitanović–Deligne exceptional series.
@@
 For instance, the Cvitanović–Deligne exceptional series displays striking uniformity in the tensor product decompositions of adjoint representations, and the dimensions of its irreducible representations admit expressions as rational functions of the dual Coxeter number Deligne (1996); Cohen and de Man (1996); Deligne and de Man (1996).
-From the perspective of Vogel’s theory of the universal Lie algebra, both the Cvitanović–Deligne series and the subexceptional series correspond to straight lines in Vogel’s projective plane Vogel (1999).
+This series appears frequently in theoretical physics, including 2d rational CFTs Mathur et al.
+(1988), 4d $N=2$ superconformal field theories (SCFTs) (see e.g.
+(Beem and Rastelli, 2018, Section 4)), and 6d $N=(1,0)$ SCFTs Morrison and Taylor (2012).
+The subexceptional series first appeared as global symmetries in 4d extended supergravity theories Cvitanovic (1981), and it also possesses many distinguished representations with uniform tensor product decompositions and dimension formulas Landsberg and Manivel (2004).
+From the perspective of Vogel’s theory of the universal Lie algebra, both the Cvitanović–Deligne series and the subexceptional series form straight lines in Vogel’s projective plane Vogel (1999).
 Moreover, the four exceptional series are deeply connected with geometry Landsberg and Manivel (2002, 2004, 2001).
@@
 These absent elements were identified as discrete groups in Deligne and Gross (2002).
-A second issue concerns the common practice in the physics literature of adjoining an additional entry “Lee–Yang model” at the beginning of the Cvitanović–Deligne exceptional series, which further obscures the triangular structure.
+A second issue concerns the common practice in the physics literature of adjoining an additional entry “Lee–Yang model” at the beginning of the Cvitanović–Deligne exceptional series Mathur et al.
+(1988); Beem and Rastelli (2018), which further obscures the triangular structure.
 A third question is whether the famous intermediate Lie algebra $E_{7+1/2}$ Landsberg and Manivel (2006), lying between $E_{7}$ and $E_{8}$, can be incorporated into the Cvitanović–Deligne exceptional series in a manner compatible with the magic triangle.
@@
 The remaining entries consist of various minimal models, compact boson theories, and their non-diagonal modular invariants, which will be explained in detail below.
-In particular, we introduce an additional row ending at $E_{7+1/2}$, referred to as the intermediate exceptional series, which has been studied recently in Lee et al.
+In particular, we introduce an additional row ending at $E_{7+1/2}$, referred to as the intermediate exceptional series, which has been studied systematically in Lee et al.
 (2024b).
@@
 1.
-We determine the 2d RCFT associated with each entry of the extended magic triangle at level one.
+We determine the 2d RCFT associated with each entry of the extended magic triangle at level one, as presented in Table 2.
 The characters of these RCFTs satisfy a uniform two-parameter family of fourth-order holomorphic modular linear differential equations (MLDEs), given in (41).
-2. The magic coset (48) provides a natural generalization of the dual pair structure with respect to $(E_{8})_{1}$ in the Cvitanović–Deligne exceptional series to the entire magic triangle.
-3. The dimensions, degeneracies, and modular $S$-matrices of all theories in the magic triangle are given uniformly.
-4.
-We show that all thirty theories on the level-one magic triangle can be decomposed into five elementary building blocks ${\rm Vir}^{\rm eff}_{5,2}$, ${\rm Vir}^{\rm eff}_{5,3}$, $(U_{1})_{3}$, ${\rm Vir}^{\rm e}_{6,5}$ and $D_{\rm 2A}$.
-5. The subexceptional series at level two exhibits $\mathcal{N}=1$ supersymmetry. The corresponding Neveu–Schwarz sector characters satisfy a one-parameter family of fourth-order fermionic MLDEs.
-6. Uniform coset constructions for $(\mathfrak{g}_{\rm CD})_{2}/{(\mathfrak{g}_{\rm sub})_{2}\times(A_{1})_{2}}$ and $(\mathfrak{g}_{\rm sub})_{2}/{(\mathfrak{g}_{\rm Severi})_{2}\times(U_{1})_{6}}$.
+2. The magic coset (48) provides a natural generalization of the dual-pair structure with respect to $(E_{8})_{1}$ in the Cvitanović–Deligne exceptional series to the entire magic triangle.
+3. The dimensions, degeneracies, and modular $S$-matrices of all theories in the magic triangle at level one are given in a uniform manner.
+4. We show that all thirty theories on the level-one magic triangle can be decomposed into five elementary building blocks: Vir^eff_5,2, Vir^eff_5,3, (U_1)_3, Vir^e_6,5, D_2A.
+5. The subexceptional series at level two exhibits N=1{N}=1 supersymmetry. The corresponding fermionic characters satisfy a one-parameter family of fourth-order fermionic MLDEs.
+6.
+We identify many uniform coset constructions involving exceptional series at level two such as $(\mathfrak{g}_{\rm CD})_{2}/{(\mathfrak{g}_{\rm sub})_{2}\times(A_{1})_{2}}$ and $(\mathfrak{g}_{\rm sub})_{2}/{(\mathfrak{g}_{\rm Severi})_{2}\times(U_{1})_{6}}$.
 This paper is organized as follows.
@@
 In Section 3, we extend the framework of Mathur–Mukhi–Sen to the full magic triangle and investigate several structural properties of this generalization, which may be regarded as an affine realization of the magic triangle at level one.
-Section 4 is devoted to the magic triangle at level two, with particular emphasis on the exceptional series appearing in the lower rows which exhibits more symmetries.
-In Section 5, we discuss the magic triangle at arbitrary higher levels.
+Section 4 is devoted to the magic triangle at level two, with particular emphasis on the exceptional series appearing in the lower rows which exhibits richer symmetry structures.
+In Section 5, we discuss the magic triangle at arbitrary positive integer levels.
 In the end, we give some concluding remarks.
-In Appendix A, we explicitly present the data for the extended magic triangle and CFTs at level one.
-In Appendix B, we explicitly give the S-matrices for the CFTs at level one of CD exceptional series.
-In Appendix C, we discuss the spurious solutions of the 4th order MLDE for CD exceptional series.
-We use $()^{\rm e}$ for a non-diagonal modular invariant of RCFT, i.e.
-a VOA extension.
+In Appendix A, we explicitly present the data for the extended magic triangle and the CFTs at level one.
+In Appendix B, we explicitly give the $S$-matrices for the CFTs at level one of Cvitanović–Deligne exceptional series.
+In Appendix C, we discuss the spurious solutions of the 4th order MLDE for Cvitanović–Deligne exceptional series.
+We use $()^{\rm e}$ for a non-diagonal modular invariant of RCFT, that is, a VOA extension.
 For example, ${\rm Vir}^{\rm e}_{6,5}$ is a non-diagonal modular invariant of Virasoro minimal model ${\rm Vir}_{6,5}$ that describes the critical three-state Potts model.
-We denote the RCFT associated to the triangle element $(\mu,\nu)$ at level $k$ as $\mathcal{T}^{(\mu,\nu)}_{k}$.
-Especially we denote the level-one theory as $\mathcal{T}^{(\mu,\nu)}$ for simplicity.
-We use subscript to denote the degeneracy of primary fields.
-We use $\rm eff$ to denote the effective theory of non-unitary minimal models.
-2 Review on the magic triangle
+We denote the RCFT associated with the triangle element $(\mu,\nu)$ at level $k$ as $\mathcal{T}^{(\mu,\nu)}_{k}$.
+In particular, we denote the level-one theory as $\mathcal{T}^{(\mu,\nu)}$ for simplicity.
+We use subscript numbers to denote the degeneracies of primary fields, and we use the superscript $\rm eff$” to denote the effective theory of non-unitary minimal models.
+2 Review of the magic triangle
 2.1 Exceptional series and magic triangle
@@
 $\{1/5,1/4,1/3,1/2,2/3,1,3/2,2,3,4,5\}.$ (1)
-They correspond to the following extended Cvitanović–Deligne exceptional series respectively
+These values correspond to the following extended Cvitanović–Deligne exceptional series, respectively:
 $e\subset A_{1/2}\subset A_{1}\subset A_{2}\subset G_{2}\subset D_{4}\subset F_{4}\subset E_{6}\subset E_{7}\subset E_{7+1/2}\subset E_{8}.$ (2)
-Here $e$ is a trivial element.
-For the simple Lie algebras $\mathfrak{g}$ in the series, which constitute the original Cvitanović–Deligne exceptional series, the dimension and the dual Coxeter number Cvitanovic (1981, 2008); Deligne (1996) are given by $\nu$ as
+Here $e$ denotes a trivial element.
+For the simple Lie algebras $\mathfrak{g}$ in the series, which constitute the original Cvitanović–Deligne exceptional series, the dimension and the dual Coxeter number Cvitanovic (1981, 2008); Deligne (1996) are expressed in terms of $\nu$ as
 $d_{\mathfrak{g}}$ $=\frac{2(5\nu-1)(6\nu+1)}{(\nu+1)},$ (3)
@@
 Moreover, infinitely many representations of $\mathfrak{g}$ admit dimension formulas that can be expressed as rational functions of $\nu$ Cohen and de Man (1996); Landsberg and Manivel (2006).
-The original CD exceptional series is extended to include intermediate Lie algebras $A_{1/2}$ and $E_{7+1/2}$, whose dual Coxeter number and dimension obey the same formula.
+The original Cvitanović–Deligne exceptional series can be extended to include intermediate Lie algebras $A_{1/2}$ Shtepin (1993) and $E_{7+1/2}$ Landsberg and Manivel (2006), whose dual Coxeter number and dimension obey the same formulas (3) and (4).
+[^1] The notation “half” indicates that the dual Coxeter number of the intermediate Lie algebra is the arithmetic mean of those of the two algebras appearing in the corresponding inclusions, namely $A_{0}\subset A_{1/2}\subset A_{1}$ and $E_{7}\subset E_{7+1/2}\subset E_{8}$.
 To be specific, $A_{1/2}$ has dual Coxeter number $3/2$ and dimension $1$.
-In many senses, one can regard $A_{1/2}$ as Lie superalgebra $osp(1|2)$ which has dual Coxeter number $3/2$ and superdimension $1$.
+In many senses, one may regard $A_{1/2}$ as Lie superalgebra $osp(1|2)$ which has dual Coxeter number $3/2$ and superdimension $1$.
 Then affine characters of $(A_{1/2})_{k}$ are just the supercharacters of $osp(1|2)_{k}$.
-An alternative interpretation is to regard $A_{1/2}$ as a formal object, denoted $LY$, such that the associated theory $(LY)k$ realizes the Lee–Yang model at level $k$, namely the effective Virasoro minimal model $\mathrm{Vir}^{\rm eff}_{(2k+3,2)}$ Duan et al.
+An alternative interpretation is to regard $A_{1/2}$ as a formal object, denoted $LY$, such that the associated theory $(LY)_{k}$ realizes the Lee–Yang model at level $k$, namely the effective Virasoro minimal model $\mathrm{Vir}^{\rm eff}_{(2k+3,2)}$ Duan et al.
 (2022).
-This correspondence follows from the fact that the supercharacters of $\mathfrak{osp}(1|2)_{k}$ agree, as $q$-series, with the characters of $\mathrm{Vir}^{\rm eff}_{(2k+3,2)}$.
-The $E_{7+1/2}$ is a famous intermediate Lie algebra with dual Coxeter number $24$ and dimension $190$.
+This correspondence follows from the fact that the supercharacters of $\mathfrak{osp}(1|2)_{k}$ agree, as $q$-series, with the characters of $\mathrm{Vir}^{\rm eff}_{(2k+3,2)}$ Ferrari et al.
+(2024).
+On the other hand, the $E_{7+1/2}$ is a famous intermediate Lie algebra with dual Coxeter number $24$ and dimension $190$.
 Its existence was noted both in physics Mathur et al.
 (1988); Cvitanovic (2008) and mathematics Cohen and de Man (1996), and it was subsequently constructed by Landsberg and Manivel using sextonions Landsberg and Manivel (2006).
+Many irreducible representations of $E_{7+1/2}$ have been studied in Lee et al.
+(2024a).
+The affine characters of $(E_{7+1/2})_{1}$ were found both by MLDE Mathur et al.
+(1988) and Hecke operator Harvey and Wu (2018).
+The affine characters of $(E_{7+1/2})_{2}$ were recently found by Hecke operator in Duan et al.
+(2022).
 The magic triangle is a two-dimensional symmetric generalization of the Cvitanović–Deligne exceptional series, parameterized by a pair $(\mu,\nu)$ with $\mu\nu\geq 1$.
-In the case $\mu\nu=1$, all entries reduce to trivial elements.
-One systematic construction of the magic triangle proceeds by placing the Cvitanovi’c–Deligne exceptional series along the bottom row and the rightmost column, and then defining each remaining entry as the centralizer of the corresponding row element inside the column element Deligne and Gross (2002).
+In the case $\mu\nu=1$, all entries reduce to trivial elements due to the zero dimension.
+[^2] They may nevertheless be interesting, and we hope to return to this point elsewhere.
+One systematic construction of the magic triangle proceeds by placing the Cvitanović–Deligne exceptional series along the bottom row and the rightmost column, and then defining each remaining entry as the centralizer of the corresponding row element inside the column element Deligne and Gross (2002).
 The original magic triangle introduced in (Cvitanovic, 1981, Table 6) is reproduced in Table 1.
 The subarray with $\mu,\nu>1$, highlighted in Table 1, is known as the Freudenthal–Tits magic square.
-The Lie algebras appearing in the magic square admit a uniform construction in terms of pairs of normed division algebras, $\mathfrak{m}(\mathbb{A}_{1},\mathbb{A}_{2})$ for $\mathbb{A}=\mathbb{R},\mathbb{C},\mathbb{H},\mathbb{O}$.
+The Lie algebras appearing in the magic square admit a uniform construction in terms of pairs of normed division algebras, denoted by $\mathfrak{m}(\mathbb{A}_{1},\mathbb{A}_{2})$ for $\mathbb{A}=\mathbb{R},\mathbb{C},\mathbb{H},\mathbb{O}$.
 This construction realizes the idea that all five exceptional Lie algebras are related to the octonions; see Baez (2002) for a comprehensive review.
-Upon including the sextonion algebra $\mathbb{S}$, the magic square can be further extended to incorporate several intermediate Lie algebras; we refer to Lee et al.
+Upon including the sextonion algebra $\mathbb{S}$, the magic square can be further extended to incorporate several intermediate Lie algebras; see Westbury (2006); Lee et al.
 (2024b) for a detailed discussion.
-In this context, it is sometimes convenient to replace the parameter $\nu$ by $a=2(\nu-1)$, which takes values in the set $\{-8/5,-3/2,-4/3,-1,-2/3,0,1,2,4,6,8\}.$ Here the numbers $1,2,4,6,8$ are the real dimension of $\mathbb{R},\mathbb{C},\mathbb{H},\mathbb{S},\mathbb{O}$.
+In this context, it is sometimes convenient to replace the parameter $\nu$ by $a=2(\nu-1)$, which takes values in the set $\{-8/5,-3/2,-4/3,-1,-2/3,0,1,2,4,6,8\}.$ Here the numbers $1,2,4,6$, and $8$ are the real dimensions of $\mathbb{R},\mathbb{C},\mathbb{H},\mathbb{S}$, and $\mathbb{O}$, respectively.
 Finally, it is worth noting that the magic square has also appeared in the study of supergravity theories, see for example Anastasiou et al.
 (2014).
-For the element ${\mathfrak{g}}^{(\mu,\nu)}$ in the original magic triangle in Table 1, the dimension and dual Coxeter number have the following universal formulas
+For the element ${\mathfrak{g}}^{(\mu,\nu)}$ in the original magic triangle in Table 1, the dimension and dual Coxeter number admit the following universal formulas:
 $d_{\mathfrak{g}}^{(\mu,\nu)}$ $=\frac{12(\mu\nu-1)(\mu\nu+\mu+\nu-4)}{(\mu+1)(\nu+1)},$ (5)
 $h^{\vee}_{\mathfrak{g}}{}^{(\mu,\nu)}$ $=\mu\nu+\mu+\nu-5.$ (6)
-The dimension formula coincides with that of the magic square derived in Landsberg and Manivel (2006).
-In particular, by specializing to $\mu=5$, one recovers the expressions for the dimensions and dual Coxeter numbers of the Cvitanović–Deligne exceptional series.
-The original magic triangle of Cvitanović and of Deligne–Gross does not include the cases $\mu,\nu=1/4$ or $4$.
-Incorporating these values leads naturally to an extended magic triangle.
-Remarkably, the universal formulas (5) and (6) continue to hold throughout this extension.
+The dimension formula was given in (Cvitanovic, 1981, Equation (11.1)) in a different form, and also given for magic square in (Landsberg and Manivel, 2006, Proposition 3.2).
+In particular, by specializing to $\mu=5$, one recovers the expressions for the dimensions (3) and dual Coxeter numbers (4) of the Cvitanović–Deligne exceptional series.
+The original magic triangle of Cvitanović and Deligne–Gross does not include the cases $\mu,\nu=1/4$ or $4$.
+Incorporating these values naturally leads to an extended magic triangle.
+Interestingly, the universal formulas (5) and (6) continue to hold throughout this extension.
+In particular, the dimensions are all non-negative integers.
 For $\nu=1/4$, the extension introduces a single additional element, namely $A_{1/2}$.
@@
 In Deligne and Gross (2002), these entries are realized concretely by the finite groups $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ and $\mathbb{Z}_{3}$, respectively.
-The formal group ${\mathfrak{g}}^{(4,1/3)}$ has dimension $1$ and dual Coxeter number $2/3$.
+The element ${\mathfrak{g}}^{(4,1/3)}$ has dimension $1$ and dual Coxeter number $2/3$.
 It may be interpreted as a formal object, denoted $IM$, such that the associated theory $(IM)_{k}$ realizes the effective minimal model $\mathrm{Vir}^{\rm eff}_{(3k+2,3)}$ Lee et al.
@@
 According to the general formula (6), their dual Coxeter numbers are given by $2/3$, $1/4$, and $3/2$, respectively.
-Equivalently, the invariant bilinear forms on these Lie algebras are rescaled by factors of $3$, 88, and $2$.
+This means that the invariant bilinear forms on these Lie algebras are rescaled by factors of $3$, 88, and $2$.
 Consequently, when constructing the associated Wess–Zumino–Witten models, the levels must be rescaled by the same factors, namely $3$, 88, and $2$, respectively.
 2.2 Modular linear differential equation
-In this section, we briefly review modular linear differential equations (MLDEs) and their fermionic version, namely MLDEs associated with congruence subgroups of level two.
+In this section, we briefly review modular linear differential equations (MLDEs) and their fermionic counter parts, namely MLDEs associated with congruence subgroups of level two.
 There exist several equivalent formulations of MLDEs, including those based on ordinary derivatives, Serre derivatives, and Rankin–Cohen brackets; see, for example, Nagatomo et al.
@@
 The non-negative integer $l$ is referred to as the index of the MLDE.
-When $l=0$, the equation is called holomorphic or monic; otherwise it is meromorphic.
+When $l=0$, the equation is called holomorphic or monic; otherwise meromorphic.
 We are interested in the special situation where $\chi$ is a modular form of weight zero on certain congruence subgroup of $SL(2,\mathbb{Z})$. In particular, we focus on solutions of the form
@@
 $\frac{l}{6}=\frac{d(d-1)}{12}-\sum_{i=0}^{d-1}\alpha_{i}.$ (9)
-When these solutions are identified with the characters of a 2d RCFT, the exponents are related to the central charge $c$ and conformal weights $h_{i}$ by
+When these solutions are identified with the characters of a 2d RCFT, the exponents are related to the central charge $c$ and conformal weights $h_{i}$ of primary fields by
 $\alpha_{i}=-\frac{c}{24}+h_{i}.$
-One of the important applications of MLDE is to classify 2d RCFTs with given number of characters, or bootstrap new 2d RCFTs.
-This program requires a classification of CFT-type solutions for given pairs $(d,l)$.
+One of the important applications of MLDEs is the classification of 2d RCFTs with a given number of characters, as well as the bootstrapping of new 2d RCFTs.
+This program requires a classification of CFT-type solutions of MLDEs for given pairs $(d,l)$.
 In the next subsection, we review the classical work of Mathur, Mukhi, and Sen on MLDEs with $(d,l)=(2,0)$ Mathur et al.
@@
 It is natural to generalize the properties of the characters of 2d RCFTs to the case of 2d fermionic RCFTs.
-In the presence of fermions, one must distinguish between periodic (Ramond) and anti-periodic (Neveu–Schwarz) boundary conditions along a circle, leading to four sectors depending on the boundary conditions along the two cycles of the torus.
-Accordingly, fermionic RCFTs admit four types of characters: NS, $\widetilde{\rm NS}$, R, and $\widetilde{\rm R}$.
-These transform as vector-valued modular forms for the congruence subgroups $\Gamma_{\theta}(2)$, $\Gamma^{0}(2)$, $\Gamma_{0}(2)$, and $SL(2,\mathbb{Z})$, respectively; see, for instance, Bae et al.
+In the presence of fermions, one must distinguish between periodic (Ramond) and antiperiodic (Neveu–Schwarz) boundary conditions along a circle, leading to four sectors depending on the boundary conditions along the two cycles of the torus.
+In summary, fermionic RCFTs admit four types of characters: NS, $\widetilde{\rm NS}$, R, and $\widetilde{\rm R}$.
+They transform as vector-valued modular forms for the congruence subgroups $\Gamma_{\theta}(2)$, $\Gamma^{0}(2)$, $\Gamma_{0}(2)$, and $SL(2,\mathbb{Z})$, respectively; see, for instance, Bae et al.
 (2021a); Duan et al.
-(2023).
+(2023) for details.
 The fermionic characters of a fermionic RCFT satisfy corresponding fermionic MLDEs.
-For example, the Neveu–Schwarz characters obey an MLDE of the form (8), where the coefficients $\phi_{k}$ are holomorphic modular forms of weight $k$ for the congruence subgroup $\Gamma_{\theta}(2)$.
+For example, the NS characters obey an MLDE of the form (8), where the coefficients $\phi_{k}$ are holomorphic modular forms of weight $k$ for the congruence subgroup $\Gamma_{\theta}(2)$.
+Then the fermionic MLDEs for $\widetilde{\rm NS}$ characters and the one for R characters can be obtained by modular transformation from the one for NS characters.
 2.3 Mathur–Mukhi–Sen’s discovery
@@
 This equation is also known in the literature as the Kaneko–Zagier equation Kaneko and Zagier (1998).
-Mathur, Mukhi, and Sen showed that only ten discrete values of the parameter $\lambda$ admit CFT-type solutions.
+Mathur, Mukhi, and Sen showed that only for ten discrete values of the parameter $\lambda$ does the MLDE (10) admit CFT-type solutions satisfying with the unique vacuum condition.
 Moreover, for each such value of $\lambda$, the corresponding solutions coincide precisely with the characters of level-one Wess–Zumino–Witten models associated with the extended Cvitanović–Deligne exceptional series:
 $\mathrm{Vir}_{5,2}^{\rm eff},\ (A_{1})_{1},\ (A_{2})_{1},\ (G_{2})_{1},\ (D_{4})_{1},\ (F_{4})_{1},\ (E_{6})_{1},\ (E_{7})_{1},\ (E_{7+1/2})_{1},\ (E_{8})_{1}.$ (11)
-We refer to the sequence (11) as the Mathur–Mukhi–Sen series, and denote them as $\mathcal{T}^{(\nu,5)}$ or $\mathcal{T}^{(5,\nu)}$ theories with $\nu$ defined earlier associated to the CD exceptional series.
+We refer to the sequence (11) as the Mathur–Mukhi–Sen series, and denote these theories as $\mathcal{T}^{(\nu,5)}$ or $\mathcal{T}^{(5,\nu)}$, where $\nu$ is defined earlier and is associated with the CD exceptional series.
 This notation will be consistent with magic triangle at level one in Table 2.
 The associated CFT data are summarized in Table 3.
-Here $\mathrm{Vir}_{5,2}^{\rm eff}$ denotes the effective Lee–Yang model with effective central charge $2/5$.
+Here $\mathrm{Vir}_{5,2}^{\rm eff}$ denotes the effective Lee–Yang model with effective central charge 2/52/5.333This is to exchange the labeling of the vacuum and non-vacuum operator of the original Lee–Yang model with c=22/5.c=-{22}/{5}.
 The entry $(E_{7+1/2})_{1}$ corresponds to the solution of the MLDE with central charge $c=38/5$.
-Mathur, Mukhi, and Sen observed that, if interpreted as a level-one WZW model, the corresponding symmetry algebra would have dimension $190$ and dual Coxeter number $24$, and hence could not arise from a semisimple Lie algebra.
+Mathur, Mukhi, and Sen observed that, if this solution were interpreted as a level-one WZW model, the corresponding symmetry algebra would have dimension $190$ and dual Coxeter number $24$, and hence could not arise from a semisimple Lie algebra.
 It was subsequently shown that these solutions can be realized as the characters of an intermediate vertex subalgebra of $(E_{8})_{1}$ Kawasetsu (2014).
-An alternative interpretation identifies them as the characters of a minimal $\mathcal{W}$-algebra with central charge $32/5$ Kawasetsu (2018).
-The $(E_{8})_{1}$ theory has the single vacuum character $j(\tau)^{1/3}$ Mathur et al.
+Accordingly, $(E_{7+1/2})_{1}$ is often reffered to as an intermediate VOA in the physics literature.
+An alternative interpretation identifies the solutions as the characters of Ramond-twisted modules of a minimal $\mathcal{W}$-algebra of type $E_{8}$ with central charge $32/5$ Kawasetsu (2018).
+Finally, the $(E_{8})_{1}$ theory has the single character, giveb by $j(\tau)^{1/3}$ Mathur et al.
 (1988).
-In this case, the second order MLDE has a spurious solution, see the Appendix C.
+In this case, the second order MLDE admits a spurious solution, see Appendix C.
 $\nu$ $1/4$ $1/3$ $1/2$ $2/3$ 11 $3/2$ 2 3 4 5
@@
 $R$ 1 2 3 7 8 26 27 56 57 -
-Table 3: 2d CFTs associated to the extended CD exceptional series at level 11.
-Recall that the Wess–Zumino–Witten model $(\mathfrak{g})k$ has central charge
-$c_{k}=\frac{kd_{\mathfrak{g}}}{k+h^{\vee}}.$ (12)
-For the extended Cvitanovi’c–Deligne exceptional series, both the dimension $d_{\mathfrak{g}}$ and the dual Coxeter number $h^{\vee}$ can be expressed as functions of the parameter $\nu$.
+Table 3: 2d CFTs associated to the extended Cvitanović–Deligne exceptional series at level 11.
+Recall that the Wess–Zumino–Witten model $(\mathfrak{g})_{k}$ has central charge and conformal weights given by
+$c_{k}=\frac{kd_{\mathfrak{g}}}{k+h^{\vee}},\qquad h_{\lambda}=\frac{C_{2}(R_{\lambda})}{2(k+h^{\vee})},$ (12)
+where $\lambda$ is the weight and $C_{2}(R_{\lambda})$ is the quadratic Casimir invariant of representation $R_{\lambda}$.
+For the extended Cvitanović–Deligne exceptional series, both the dimension $d_{\mathfrak{g}}$ and the dual Coxeter number $h^{\vee}$ can be expressed as functions of the parameter $\nu$; the same is true for the quadratic Casimir invariant of the fundamental representation.
 At level one, the corresponding central charge and the conformal weight of the non-vacuum primary field are given by
 $c=\frac{2(5\nu-1)}{(\nu+1)},\qquad h=\frac{\nu}{\nu+1}.$ (13)
-In terms of the parameter $\nu$, the associated second-order holomorphic MLDE can be written compactly as $\Delta_{\nu}\chi=0$, where $\Delta_{\nu}\chi=0$, where
+In terms of the parameter $\nu$, the associated second-order holomorphic MLDE can be written compactly as $\Delta_{\nu}\chi=0$, where
 $\Delta_{\nu}=D^{2}-\frac{(5\nu-1)(7\nu+1)}{144(\nu+1)^{2}}E_{4}.$ (14)
@@
 $Z=|\chi_{0}|^{2}+{\mathcal{N}}|\chi_{1}|^{2},$ (15)
-where $\mathcal{N}$ is natural integer, capturing the degeneracy of the non-vaccum primary operator. The characters of the Mathur–Mukhi–Sen series admit the expansions
-$\chi_{0}$ $=q^{-c/24}(1+d_{\mathfrak{g}}q+\cdots),$
-$\chi_{1}$ $=q^{-c/24+h}(R+\cdots),$ (16)
-where $R$ denotes the dimension of the fundamental representation of $\mathfrak{g}$. The $S$-matrix arises from the modular transformation of the characters,
+where $\mathcal{N}$ is a natural number, capturing the degeneracy of the non-vacuum primary operator. The characters of the Mathur–Mukhi–Sen series admit the Fourier expansions
+$\chi_{0}$ $=q^{-\frac{c}{24}}(1+d_{\mathfrak{g}}q+\cdots),$
+$\chi_{1}$ $=q^{-\frac{c}{24}+h}(R+\cdots),$ (16)
+where $R$ denotes the dimension of the fundamental representation of $\mathfrak{g}$.
+The degeneracy $\mathcal{N}$ and dimension $R$ for various $\nu$ are collected in Table 3.
+The $S$-matrix arises from the modular transformation of the characters,
 $\left(\begin{array}[]{c}\chi_{0}\\ \chi_{1}\end{array}\right)(-1/\tau)=S\cdot\left(\begin{array}[]{c}\chi_{0}\\ \chi_{1}\end{array}\right)(\tau).$ (21)
-As a $2\times 2$ matrix, the $S$-matrix satisfies $\det S_{(\nu)}=-1$ and
+As a $2\times 2$ matrix not necessarily symmetric, the $S$-matrix satisfies $\det S_{(\nu)}=-1$ and
 $S_{(\nu)}^{T}\cdot\left(\begin{array}[]{cc}1&0\\ 0&\mathcal{N}_{(\nu)}\end{array}\right)\cdot S_{(\nu)}=\left(\begin{array}[]{cc}1&0\\ 0&\mathcal{N}_{(\nu)}\end{array}\right).$ (26)
@@
 More generally, one may consider the full modular $S_{(\nu)}$-matrix as a $(1+\mathcal{N}_{(\nu)})$-dimensional unitary matrix.
-It is well known that for the CD exceptional series at level $1$, the RCFTs pair together as a coset with respect to $\mathcal{T}^{(5,5)}=(E_{8})_{1}$ Mathur et al. (1988). These cosets are
+It is well known that for the CD exceptional series at level $1$, the RCFTs pair together as a coset with respect to $\mathcal{T}^{(5,5)}=(E_{8})_{1}$ Mathur et al. (1988). These coset relations are
 $(E_{7})_{1}=\frac{(E_{8})_{1}}{(A_{1})_{1}},\quad(E_{6})_{1}=\frac{(E_{8})_{1}}{(A_{2})_{1}},\quad(F_{4})_{1}=\frac{(E_{8})_{1}}{(G_{2})_{1}},\quad(D_{4})_{1}=\frac{(E_{8})_{1}}{(D_{4})_{1}},\quad$ (27)
 In our notation, these cosets can be written uniformly as
-$\mathcal{T}^{(5,\nu)}=\frac{(E_{8})_{1}}{\mathcal{T}^{(1/\nu,5)}}.$ (28)
-The non-vacuum conformal weight $h_{\nu}$ of $\mathcal{T}_{5,\nu}$ satisfies the following relation $h_{\nu}+h_{1/\nu}=1.$ The degeneracy $\mathcal{N}_{(\nu)}$ and $S$-matrix $S_{(\nu)}$ satisfy
-$\mathcal{N}_{(\nu)}=\mathcal{N}_{(1/\nu)},\ \ S_{(\nu)}=S_{(1/\nu)}.$ (29)
+$\mathcal{T}^{(5,\nu)}=\frac{\mathcal{T}^{(5,5)}}{\mathcal{T}^{(1/\nu,5)}}.$ (28)
+The non-vacuum conformal weight $h_{\nu}$ of $\mathcal{T}_{5,\nu}$ theory satisfies the relation $h_{\nu}+h_{1/\nu}=1.$ The degeneracy $\mathcal{N}_{(\nu)}$ and $S$-matrix $S_{(\nu)}$ satisfy
+$\mathcal{N}_{(\nu)}=\mathcal{N}_{(1/\nu)},\qquad S_{(\nu)}=S_{(1/\nu)}.$ (29)
 The characters satisfy the relation
-$\chi_{0}^{(\nu)}\chi_{0}^{(1/\nu)}+\mathcal{N}_{(\nu)}\chi_{1}^{(\nu)}\chi_{1}^{(1/\nu)}=j^{1/3},$ (30)
-which is $S$-invariant. The above bilinear relation in the subleading $q$-order implies
+$\chi_{0}^{(\nu)}\chi_{0}^{(1/\nu)}+\mathcal{N}_{(\nu)}\chi_{1}^{(\nu)}\chi_{1}^{(1/\nu)}=j^{1/3}.$ (30)
+The above bilinear relation of characters in the subleading $q$-order implies
 $\mathcal{N}_{(\nu)}R_{(\nu)}R_{(1/\nu)}=312-60(\nu+1/\nu).$ (31)
@@
 $\mathcal{N}_{(\nu)}R_{(\nu)}^{2}=2^{16h}(4\sin^{2}(\pi h)-1)\bigg[\frac{\Gamma(1-2h)\Gamma(1/2+3h)}{\Gamma(1+2h)\Gamma(1/2-h)}\bigg]^{2}.$ (32)
-Multiplying $\mathcal{N}_{(\nu)}R_{(1/\nu)}^{2}$ one can also obtain the relation (31).
-The characters of WZW models for CD exceptional series at level one have many other interesting properties such as Hecke relations Harvey and Wu (2018).
-Besides, it is recently shown that the vacuum character of WZW models for CD exceptional series at level 1 can be constructed from the monodromy trace of some Argyres–Douglas theories Kim and Song (2024).
+By multiplying $\mathcal{N}_{(\nu)}R_{(1/\nu)}^{2}$ one can also obtain the relation (31).
+The characters of WZW models for the CD exceptional series at level one possess many other interesting properties such as Hecke relations Harvey and Wu (2018).
+Besides, it is recently shown that the vacuum character of WZW models for CD exceptional series at level one can be constructed from the monodromy trace of some Argyres–Douglas theories Kim and Song (2024).
 3 Magic triangle at level one
-A principal motivation of this work is to extend the discovery of Mathur, Mukhi, and Sen concerning second-order holomorphic MLDEs and the Cvitanovi’c–Deligne exceptional series to the full magic triangle.
+A principal motivation of this work is to extend the discovery of Mathur, Mukhi, and Sen concerning second-order holomorphic MLDEs and the Cvitanović–Deligne exceptional series to the full magic triangle.
 To this end, we begin by analyzing the level-one Wess–Zumino–Witten models associated with the ordinary Lie groups appearing in the original magic triangle of Table 1.
-For each resulting rational conformal field theory, we compute the characters and show that they satisfy a uniform two-parameter family of fourth-order holomorphic MLDEs.
-Assuming that this MLDE framework extends to the enlarged magic triangle, we solve for the characters corresponding to the previously unknown entries and identify the associated 2d RCFTs.
+For each resulting rational conformal field theory, we compute the characters and find that they satisfy a uniform two-parameter family of fourth-order holomorphic MLDE (41).
+Assuming that this MLDE framework extends to the enlarged magic triangle, we determine the characters corresponding to the previously unknown entries and identify the associated 2d RCFTs.
 The complete set of resulting theories is summarized in Table 2, which we interpret as the extended magic triangle at level one.
-We further uncover several structural properties of this level-one triangle, including the magic coset, magic bilinear relation and matrix-valued modularity of the characters.
+We further uncover several elegant properties of this level-one triangle, including the magic coset structure, matrix-valued modularity of the characters with explicit S-matrices, the dimensions and degeneracies of primary operators, and the existence of five atomic models that together constitute all theories in the triangle.
 3.1 The fourth order MLDE
 Let us first consider the level-one theories for the original magic triangle in Table 1.
-Substituting the dimension and dual Coxeter number formulas (5), (6) into the central charge formula of WZW model (12), we obtain the central charge for the level-one theory $\mathcal{T}^{(\mu,\nu)}$ as
+Substituting the dimension and dual Coxeter number formulas (5), (6) into the central charge formula of WZW model (12), we obtain the central charge of the level-one theory $\mathcal{T}^{(\mu,\nu)}$ as
 $c^{(\mu,\nu)}=\frac{12(\mu\nu-1)}{(\mu+1)(\nu+1)}.$ (33)
@@
 $h_{i}^{(\mu,\nu)}=0,\ \frac{\mu\nu-1}{(\mu+1)(\nu+1)},\ \frac{\nu}{\nu+1},\ \frac{\mu}{\mu+1}.$ (34)
-The exponents $\alpha_{i}=-c/24+h_{i}$ in the corresponding characters $\chi_{i}$ take values,
+The exponents $\alpha_{i}=-c/24+h_{i}$, appearing in the corresponding characters $\chi_{i}$, take values
 $\alpha_{0}=-\alpha_{1}=\frac{1-\mu\nu}{2(\mu+1)(\nu+1)},\quad\alpha_{2}=\frac{\mu\nu+2\nu+1}{2(\mu+1)(\nu+1)},\quad\alpha_{3}=\frac{\mu\nu+2\mu+1}{2(\mu+1)(\nu+1)}.$ (35)
-Denote the characters of the theory $\mathcal{T}^{(\mu,\nu)}$ as
+We denote the four characters of the theory $\mathcal{T}^{(\mu,\nu)}$ by
 $\chi_{0}$ $=q^{\alpha_{0}}\Big(1+\sum_{n=1}^{\infty}a_{n}q^{n}\Big)\ \ {\rm with}\ a_{1}=d_{\mathfrak{g}},$ (36)
@@
 $\chi_{3}$ $=q^{\alpha_{3}}\Big(R_{3}+\sum_{n=1}^{\infty}d_{n}q^{n}\Big),$ (39)
-where $R_{i}$ are the dimensions of the irreducible representation of $\mathfrak{g}^{(\mu,\nu)}$.
-The coefficients $a_{n},b_{n},c_{n},d_{n}$ are the sum of the dimensions of the irreducible representations for the Lie algebra.
+where $R_{i}$ denote the dimensions of irreducible representations of $\mathfrak{g}^{(\mu,\nu)}$.
+The coefficients $a_{n},b_{n},c_{n}$, and $d_{n}$ are the sum of the dimensions of the irreducible representations for the Lie algebra.
 The modular invariant partition function of $\mathcal{T}^{(\mu,\nu)}$ on torus is
 $Z=|\chi_{0}|^{2}+\sum_{i=1}^{3}\mathcal{M}_{i}|\chi_{i}|^{2},$ (40)
-where $\mathcal{M}_{i}$ are natural numbers that capture the degeneracy of primaries fields.
-The first main result of this paper is as follows.
-We find that the characters $\chi_{i}$ of the theory $\mathcal{T}^{(\mu,\nu)}$ satisfy the following uniform fourth-order modular linear differential equation
+where $\mathcal{M}_{i}$ are positive integers that capture the degeneracies of primary fields.
+The first main result of this paper is the following.
+We find that the characters $\chi_{i}$ of the theory $\mathcal{T}^{(\mu,\nu)}$ satisfy the uniform fourth-order modular linear differential equation
 $[D^{4}+\lambda_{1}E_{4}D^{2}+\lambda_{2}E_{6}D+\lambda_{3}E_{4}^{2}]\chi=0,$ (41)
-where the three coefficients $\lambda_{1},\lambda_{2}$ and $\lambda_{3}$ are characterized by $(\mu,\nu)$ as follows:
+where the three coefficients $\lambda_{1},\lambda_{2}$ and $\lambda_{3}$ are determined by $(\mu,\nu)$ as follows:
 $\lambda_{1}$ $=-\frac{11\mu^{2}\nu^{2}+4\mu^{2}\nu+11\mu^{2}+4\mu\nu^{2}-28\mu\nu+4\mu+11\nu^{2}+4\nu+11}{36(\mu+1)^{2}(\nu+1)^{2}},$ (42)
@@
 $\lambda_{3}$ $=-\frac{(\mu\nu-1)^{2}(\mu\nu+2\mu+1)(\mu\nu+2\nu+1)}{16(\mu+1)^{4}(\nu+1)^{4}}.$ (44)
-Taking $\nu=5$, the 4th order MLDE factorizes to
+Specializing to $\nu=5$, the 4th order MLDE factorizes as
 $\left(D^{2}-\frac{(\mu-5)(\mu+7)}{144(\mu+1)^{2}}E_{4}\right)\left(D^{2}-\frac{(5\mu-1)(7\mu+1)}{144(\mu+1)^{2}}E_{4}\right)\chi=0.$ (45)
-The differential operator can be written as $\Delta_{1/\mu}\circ\Delta_{\mu}$ where $\Delta_{\mu}$ is the second order operator given in (14).
+This differential operator can thus be written as $\Delta_{1/\mu}\circ\Delta_{\mu}$ where $\Delta_{\mu}$ is the second order operator given in (14).
 This factorization demonstrates that the fourth-order MLDE (41) provides a natural generalization of the second-order Mathur–Mukhi–Sen equation.
@@
 This assumption allows us to identify two-dimensional rational conformal field theories beyond the original magic triangle of ordinary Lie groups.
-Indeed, upon specializing to $\nu=4$, the equation (41) reduces to a one-parameter family of fourth-order MLDEs governing minimal $\mathcal{W}$-algebras Kawasetsu and Sakai (2018).
-In this case, the characters of the theory $\mathcal{T}^{(\mu,4)}$ coincide with the characters of Ramond-twisted irreducible modules of the minimal $\mathcal{W}$-algebra $\mathcal{W}{-h^{\vee}/6}(\mathfrak{g},f\theta)$, where $\mathfrak{g}$ belongs to the Cvitanovi’c–Deligne exceptional series.
+Indeed, upon specializing to $\nu=4$, we find the equation (41) reduces to a one-parameter family of fourth-order MLDEs governing some minimal $\mathcal{W}$-algebras Kawasetsu and Sakai (2018).
+In this case, the characters of the theory $\mathcal{T}^{(\mu,4)}$ coincide with the characters of Ramond-twisted irreducible modules of the minimal $\mathcal{W}$-algebra $\mathcal{W}_{-h^{\vee}/6}(\mathfrak{g},f_{\theta})$, where $\mathfrak{g}$ belongs to the Cvitanović–Deligne exceptional series.
 More generally, by explicitly solving the MLDE (41) and interpreting its solutions as RCFT characters, we are led to identify a class of theories $\mathcal{T}^{(\mu,\nu)}$ extending those associated with the original magic triangle.
@@
 (2002), see Remark 5.8 there for all its irreducible modules.
-The RCFT is denoted as $M_{2}$ in (Duan et al., 2022, Section 5.6).
-The four characters have appeared as CFT-type solutions of fourth order holomorphic MLDE in (Kaidi et al., 2021, Table 6).
-The $q$-series of the characters can be found in (Duan et al., 2022, Equation (5.34)).
+This RCFT is denoted as $M_{2}$ in (Duan et al., 2022, Section 5.6).
+The four characters have appeared as CFT-type solutions of fourth order holomorphic MLDE in (Kaidi et al., 2021, Table 6), and their $q$-series expansions can be found in (Duan et al., 2022, Equation (5.34)).
 The $M_{2}$ can be realized by the coset construction $(D_{4})_{1}/(A_{2})_{1}$ (Duan et al., 2022, Section 5.6).
@@
 It has $c=6/5$ and conformal weights $h_{i}=0,(\frac{1}{10})_{3},(\frac{1}{2})_{3},\frac{3}{5}$.
-The VOA can be constructed using a code in $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ and as an extension of the product of Ising model $\mathrm{Vir}_{4,3}$ and tri-critical Ising model $\mathrm{Vir}_{5,4}$ Lam and Yamada (2000).
+The corresponding VOA can be constructed by using a code over $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ and as an extension of the product of Ising model $\mathrm{Vir}_{4,3}$ and tri-critical Ising model $\mathrm{Vir}_{5,4}$ Lam and Yamada (2000).
 For 2d RCFTs associated to various conjugacy classes and sporadic groups, see a good review in Bae et al.
 (2021b).
-  • $\mathcal{T}^{(3,2/3)}=\mathrm{Vir}_{6,5}^{\rm e}$ is a non-diagonal modular invariant of minimal model $\mathrm{Vir}_{6,5}$ which describes the 3-state Potts model.
-The VOA can be constructed using a code in $\mathbb{Z}_{3}$ Miyamoto (2001).
-  • $\mathcal{T}^{(3/2,3/2)}=(A_{1})_{8}^{\rm e}$ is the D-type modular invariant of $(A_{1})_{8}$. The VOA can be constructed by simple current extension.
-All RCFTs associated to the extended magic triangle are presented in Table 2. The data of the extended magic triangle and the RCFTs are collected in the Tables 5, 6, 7 in the Appendix A.
-Now we discuss the symmetry of the magic triangle at level one.
-It is easy to verify that under the exchange of $\mu$ and $\nu$, the central charge $c$, the conformal weight $h_{1}$ and the fourth order MLDE are invariant, while $h_{2},h_{3}$ are exchanged.
-This suggests that the $\mathcal{T}^{(\mu,\nu)}$ and $\mathcal{T}^{(\nu,\mu)}$ theories are isomorphic to each other, with only the labeling of $\chi_{2}$ and $\chi_{3}$ is exchanged.
-Therefore we only need to present half of the triangle in the Tables 5, 6, 7.
+  • $\mathcal{T}^{(3,2/3)}=\mathrm{Vir}_{6,5}^{\rm e}$ is a non-diagonal modular invariant of minimal model $\mathrm{Vir}_{6,5}$ which describes the critical 3-state Potts model.
+The corresponding VOA can be constructed by using a code over $\mathbb{Z}_{3}$ Miyamoto (2001).
+  • $\mathcal{T}^{(3/2,3/2)}=(A_{1})_{8}^{\rm e}$ is the D-type modular invariant of $(A_{1})_{8}$. The associated VOA can be constructed by simple current extension.
+All RCFTs associated with the extended magic triangle are presented in Table 2.
+The data of the extended magic triangle and the corresponding RCFTs are collected in the Tables 5, 6, 7 in the Appendix A.
+We now discuss the symmetry of the magic triangle at level one.
+It is straightforward to verify that under the exchange of $\mu$ and $\nu$, the central charge $c$, the conformal weight $h_{1}$ and the fourth order MLDE are invariant, while $h_{2},h_{3}$ are exchanged.
+This suggests that the $\mathcal{T}^{(\mu,\nu)}$ and $\mathcal{T}^{(\nu,\mu)}$ theories are isomorphic to each other, differing only by a relabeling of $\chi_{2}$ and $\chi_{3}$.
+Consequently, it suffices to present only half of the triangle, as shown in the Tables 5, 6, 7.
 In total, there are 30 non-equivalent RCFTs in the extended magic triangle at level one.
-The 16 of them come from ordinary WZW model of semisimple Lie groups, 8 of them come from intermediate CFTs, three of them from Virasoro minimal models, two of them from $U(1)$ theory, i.e.
-compact bosons on a circle, and the last one $D_{2\rm A}$ associated to Monster group.
-For all 30 theories on the magic triangle, the Fourier coefficients of the characters remain non-negative integers.
-A few remarks are in order:
+The sixteen of these theories arise from ordinary WZW model of semisimple Lie groups, eight arise from intermediate VOAs, three arise from Virasoro minimal models, two arise from $U_{1}$ theories (that is, compact bosons on a circle), and the final theory $D_{2\rm A}$ arises from the product of Ising model and tri-critical Ising model and is associated with Monster group.
+For all 30 theories on the magic triangle at level one, the Fourier coefficients of the characters remain non-negative integers.
+Several remarks are in order:
   • We find that the four characters of each theory $\mathcal{T}^{(\mu,\nu)}$ satisfy the following remarkable identity:
@@
 3.2 Magic coset
-We find that the coset relation (28) within the Cvitanovi’c–Deligne exceptional series is in fact a very special instance of a much broader class of coset constructions that exist among all theories in the extended magic triangle.
-Remarkably, for any two theories in the triangle one may form a coset that yields a third theory, which again belongs to the same triangle.
+We find that the coset relation (28) within the Cvitanović–Deligne exceptional series is in fact a very special instance of a much broader class of coset constructions that exist among all theories in the extended magic triangle.
+Remarkably, for any two theories in the magic triangle one may form a coset that yields a third theory, which again belongs to the magic triangle.
 More precisely, we uncover the following universal coset relation valid throughout the extended magic triangle:
@@
 (2024b).
-Equation (48) indicates that by taking cosets among exceptional series one can generate every theory in the extended magic triangle at level one.
+Equation (48) also indicates that by taking cosets among exceptional series one can generate every theory in the extended magic triangle at level one.
 An equivalent formulation of the magic coset relation (48) is
@@
 $\mathcal{T}^{(\mu,\nu)}=\frac{\mathcal{T}^{(\mu,5)}}{\mathcal{T}^{(1/\nu,5)}}=\frac{\mathcal{T}^{(5,\nu)}}{\mathcal{T}^{(5,1/\mu)}}.$ (52)
-This form is particularly useful, as it implies that every theory in the extended magic triangle can be realized as a coset between two theories belonging to the Cvitanović–Deligne exceptional series.
+This form is particularly useful, as it implies that every theory in the extended magic triangle in Table 2 can be realized as a coset between two theories belonging to the Cvitanović–Deligne exceptional series.
 In this sense, the magic coset construction provides a CFT analogue of the centralizer construction of Deligne–Gross for the original magic triangle Deligne and Gross (2002).
-Recall the degeneracy of non-vacuum primary of $\mathcal{T}^{(5,\nu)}$ theory is denoted as $\mathcal{N}_{(\nu)}$, which satisfies $\mathcal{N}_{(\nu)}=\mathcal{N}_{(1/\nu)}.$ The coset (52) implies that the degeneracies of $\mathcal{T}^{(\mu,\nu)}$ theory are
-$\mathcal{M}^{(\mu,\nu)}_{1,2,3}=\left(\mathcal{N}_{(\mu)}\mathcal{N}_{(\nu)},\mathcal{N}_{(\nu)},\mathcal{N}_{(\mu)}\right).$ (53)
+Recall the degeneracy of the non-vacuum primary field of $\mathcal{T}^{(5,\nu)}$ theory is denoted as $\mathcal{N}_{(\nu)}$, which satisfies $\mathcal{N}_{(\nu)}=\mathcal{N}_{(1/\nu)}.$ The numbers are collected in Table 3.
+The coset (52) implies that the degeneracies of $\mathcal{T}^{(\mu,\nu)}$ theory are
+$(\mathcal{M}^{(\mu,\nu)}_{1},\mathcal{M}^{(\mu,\nu)}_{2},\mathcal{M}^{(\mu,\nu)}_{3})=\left(\mathcal{N}_{(\mu)}\mathcal{N}_{(\nu)},\mathcal{N}_{(\nu)},\mathcal{N}_{(\mu)}\right).$ (53)
 This shows that $\mathcal{M}_{2}$ remains constant along each column, while $\mathcal{M}_{3}$ remains constant along each row of the magic triangle at level one, as summarized in Table 7.
 The magic coset (52) further determines the modular $S$-transformation of $\mathcal{T}^{(\mu,\nu)}$.
-While one might naively expect the solutions of the fourth-order MLDE to form a degree-four vector-valued modular form, it is more natural in the present context to regard the four characters as forming a $2\times 2$ matrix-valued modular form.
+While one might expect the solutions of the fourth-order MLDE to form a degree-four vector-valued modular form, it is more natural in the present context to regard the four characters as forming a $2\times 2$ matrix-valued modular form.
 [^5] We thank Don Zagier for suggesting the interpretation in terms of matrix-valued modular forms.
-Let us define the $2\times 2$ matrix $X^{(\mu,\nu)}$ for the characters $\chi_{i}$ of $\mathcal{T}^{(\mu,\nu)}$ theory as
-$X^{(\mu,\nu)}=\begin{pmatrix}\chi_{0}&\chi_{2}\\ \chi_{3}&\chi_{1}\end{pmatrix}$ (54)
+Let us define the $2\times 2$ matrix $X^{(\mu,\nu)}$ for the characters of $\mathcal{T}^{(\mu,\nu)}$ theory as
+$X^{(\mu,\nu)}=\begin{pmatrix}\chi_{0}&\chi_{2}\\ \chi_{3}&\chi_{1}\end{pmatrix}$. (54)
 We then find the following elegant $S$-transformation law:
 $X^{(\mu,\nu)}(-1/\tau)=S_{(\mu)}\cdot X^{(\mu,\nu)}(\tau)\cdot S_{(\nu)}^{T},$ (55)
-where $S_{(\nu)}$ denotes the $2\times 2$ $S$-matrix of the theory $\mathcal{T}^{(5,\nu)}$ in the Cvitanović–Deligne exceptional series.
-Taking the determinant of (55) and using $\det S_{(\nu)}=-1$, we see that $\det(X^{(\mu,\nu)})=\chi_{0}\chi_{1}-\chi_{2}\chi_{3}$ is modular invariant.
+where $S_{(\nu)}$ denotes the $2\times 2$ $S$-matrix of the theory $\mathcal{T}^{(5,\nu)}$ in the Cvitanović–Deligne exceptional series, given explicitly in B.
+Taking the determinant of (55) and using the fact $\det S_{(\nu)}=-1$, we see that $\det(X^{(\mu,\nu)})=\chi_{0}\chi_{1}-\chi_{2}\chi_{3}$ is modular invariant.
 Moreover, since $\alpha_{0}+\alpha_{1}=0$ and $\alpha_{2}+\alpha_{3}=1$, this determinant is holomorphic and must therefore be constant.
@@
 $\left(\begin{array}[]{c}\chi_{0}\\ \chi_{3}\end{array}\right)^{(\mu,5)}$ $=X^{(\mu,\nu)}\cdot\left(\begin{array}[]{cc}1&0\\ 0&\mathcal{N}_{(\nu)}\end{array}\right)\cdot\left(\begin{array}[]{c}\chi_{0}\\ \chi_{3}\end{array}\right)^{(1/\nu,5)}.$ (67)
-Extracting the leading terms yields the following relations among representation dimensions:
+Extracting the leading terms in $q$-expansion yields the following relations among representation dimensions:
 $R_{1}^{(\mu,\nu)}=\frac{\mu\nu}{12(5\mu-1)(5\nu-1)}R_{(\mu)}R_{(\nu)},$ (68)
@@
 Here $R_{(\nu)}$ denotes the dimension of the fundamental representation of the Cvitanović–Deligne exceptional series, cf.
-(2.3).
+(2.3) and Table 3.
 These relations are readily checked to be consistent with (47).
 For the row $\mathcal{T}^{(5,\nu)}$, the expression for $R_{3}^{(\mu,\nu)}$ becomes linear in $\nu$ and takes the form
-$E_{7\frac{1}{2}}:\ 12\nu-3,\ E_{7}:12\nu-4,\ E_{6}:6\nu-3,\ F_{4}:6\nu-4$ (71)
-$D_{4}:2\nu-2,\ G_{2}:2\nu-3,\ A_{2}:\nu-2,\ A_{1}:\nu-3,\ A_{\frac{1}{2}}:\nu-4.$ (72)
+$E_{7+1/2}:12\nu-3,\quad E_{7}:12\nu-4,\quad E_{6}:6\nu-3,\quad F_{4}:6\nu-4,$ (71)
+$D_{4}:2\nu-2,\quad G_{2}:2\nu-3,\quad A_{2}:\nu-2,\quad A_{1}:\nu-3,\quad A_{1/2}:\nu-4.$ (72)
 This precisely reproduces the extension of the preferred representation dimensions found by Deligne–Gross (Deligne and Gross, 2002, Equation 2).
@@
 3.3 Five atomic theories
-An interesting consequence of the magic coset relation (48) is that all thirty CFTs appearing in the extended magic triangle in Table 2 can be generated from only five atomic theories lying below the line $\mu\nu=1$.
+An interesting consequence of the magic coset relation (48) is that all thirty CFTs appearing in the extended magic triangle in Table 2 can be generated from only five atomic theories lying just below the line $\mu\nu=1$.
 These five atoms are
-$\mathrm{Vir}_{5,2}^{\rm eff},\ \mathrm{Vir}_{5,3}^{\rm eff},\ U(1)_{3},\ \mathrm{Vir}_{6,5}^{\rm e},\ D_{2\rm A}.$ (80)
-This means that every theory $\mathcal{T}^{(\mu,\nu)}$ can be realized as an extension of a product of these atomic theories, or equivalently as a non-diagonal modular invariant of an tensor product CFT.
+$\mathrm{Vir}_{5,2}^{\rm eff},\ \mathrm{Vir}_{5,3}^{\rm eff},\ (U_{1})_{3},\ \mathrm{Vir}_{6,5}^{\rm e},\ D_{2\rm A}.$ (80)
+This means that every theory $\mathcal{T}^{(\mu,\nu)}$ can be realized as an extension of a product of these atomic theories, or equivalently as a non-diagonal modular invariant of an tensor product CFT of the atoms.
 More concretely, given a theory $\mathcal{T}^{(\mu,\nu)}$ in Table 2, we draw a right triangle bounded by the three lines $\mu=\mathrm{const}$, $\nu=\mathrm{const}$, and $\mu\nu=1$.
 The theory $\mathcal{T}^{(\mu,\nu)}$ is then constructed from the atomic theories that lie along the segment of the slope just below the line $\mu\nu=1$ within this triangle.
-Since each of the five atoms appears twice along this line, any theory in the extended magic triangle involves at most two identical atomic constituents.
+Since each of the five atoms appears twice just below the $\mu\nu=1$ line, any theory in the extended magic triangle involves at most two identical atomic constituents.
 The basic data of these five atoms are summarized in Table 4.
@@
 $\mathcal{T}^{(\mu,1/\nu)}=\frac{\mathcal{T}^{(\mu,\mu)}}{\mathcal{T}^{(\mu,\nu)}},\qquad\mu>\nu.$ (81)
-This relation provides a direct generalization of the $E_{8}$ dual pairs appearing in (28).
+This relation provides a direct generalization of the $(E_{8})_{1}$ dual pairs appearing in (28).
 For instance, when $\mu=3$, equation (81) implies that the RCFTs in the subexceptional series at level one naturally pair via coset constructions with respect to $(D_{6})_{1}$.
-Given the special interest on $\mathcal{T}^{(\mu,\mu)}$ theories which are
+Given the special interest on $\mathcal{T}^{(\mu,\mu)}$ theories, namely,
 $(A_{1})_{8}^{\rm e},\ (A_{2})_{1}^{2},\ (D_{6})_{1},\ (D_{6+1/2+1/2})_{1},\ (E_{8})_{1},$
-we call them the diagonal exceptional series.
+we refer to them as the diagonal exceptional series.
 Each of these theories admits a $\mathbb{Z}_{2}$ inner automorphism such that $h_{2}=h_{3}$.
@@
 4 Magic triangle at level two
-At level two, the magic triangle no longer behaves as perfect as at level one.
-For the 18 elements belonging to the original magic triangle in Table 1, the associated higher-level theories present no conceptual difficulties, and the corresponding RCFTs are well defined.
+At level two, the magic triangle no longer behaves as perfectly as at level one.
+For the 18 elements that belong to the original magic triangle in Table 1, the associated higher-level theories do not present conceptual difficulties, and the corresponding RCFTs are well defined.
 By contrast, the remaining 12 theories in the extended triangle require more careful treatment at higher levels.
@@
 On the other hand, some intermediate Lie algebras do not admit consistent higher-level generalizations.
-A typical example is the $E_{7\frac{1}{2}}$ case, for which we have argued that only the levels $k\leq 5$ can be realized as RCFTs Lee et al.
+A typical example is the case $E_{7\frac{1}{2}}$, for which we have argued that only levels $k\leq 5$ can be realized as RCFTs Lee et al.
 (2024a).
 Using the dimension and dual Coxeter number formulas of the magic triangle, one obtains a universal expression for the level-two central charge,
-$c_{2}=\frac{24(\mu\nu-1)(\mu\nu+\mu+\nu-4)}{(\mu+1)(\nu+1)(\mu\nu+\mu+\nu-3)}$ (88)
+$c_{2}=\frac{24(\mu\nu-1)(\mu\nu+\mu+\nu-4)}{(\mu+1)(\nu+1)(\mu\nu+\mu+\nu-3)}.$ (88)
 In principle, all 13 conformal weights can be written explicitly as functions of $\mu$ and $\nu$.
-In particular, there is always a vacuum character with exponent zero.
+In particular, there is always a character with exponent zero.
 We expect that the characters of all level-two theories satisfy a two-parameter family of holomorphic modular linear differential equations of order 13 with index $l=36$.
@@
 (2024a).
-Moreover, the intermediate exceptional series at level two are realized by minimal $\mathcal{W}$-algebras $\mathcal{W}_{-h^{\vee}/6+1}(\mathfrak{g}_{\rm CD},f_{\theta})$, which has been studied recently in Sun (2024); Arakawa et al.
+Moreover, the intermediate exceptional series at level two are realized by minimal $\mathcal{W}$-algebras $\mathcal{W}_{-h^{\vee}/6+1}(\mathfrak{g}_{\rm CD},f_{\theta})$, which has been studied in Arakawa and Moreau (2018); Sun (2024); Arakawa et al.
 (2024); Kac and Wakimoto (2025).
@@
 In this case, it is most natural to fermionize the theories and study their fermionic characters and fermionic MLDEs.
-A detailed analysis of this structure will be presented in the next section.
+A detailed analysis of this structure will be presented in the next subsection.
 Finally, we expect that the Severi exceptional series at level two should admit a $\mathbb{Z}_{3}$ parafermionic structure.
@@
 4.1 Subexceptional series at level two
-A remarkable phenomenon we observe is that all theories in subexceptional series $(\mu,\nu)=(3,\nu)$ at level two are supersymmetric. These theories $\mathcal{T}_{2}^{(3,\nu)}$ are
+A remarkable phenomenon we observe is that all theories in subexceptional series at level two are supersymmetric. These theories $\mathcal{T}_{2}^{(3,\nu)}$ are
 $(U_{1})_{6},\ (A_{1})_{6},\ (A_{1}^{\oplus 3})_{2},\ (C_{3})_{2},\ (A_{5})_{2},\ (D_{6})_{2},\ (D_{6+1/2})_{2},\ (E_{7})_{2}.$ (89)
@@
 It is easy to verify this is true for all theories in (88).
-In many cases, the supersymmetry of these theories has been studied, see e.g.
+In many cases, the supersymmetry of these theories has been studied extensively, see e.g.
 Johnson-Freyd (2021); Bae and Lee (2021).
-In particular, the $(D_{6+1/2})_{2}$ was found to be supersymmetric recently in Lee et al.
+In particular, the $(D_{6+1/2})_{2}$ was found to be supersymmetric only recently in Lee et al.
 (2024b).
-Taking $\mu=3$ in (88), the central charge of $\mathcal{T}_{2}^{(3,\nu)}$ theory is
+setting $\mu=3$ in (88), the central charge of $\mathcal{T}_{2}^{(3,\nu)}$ theory becomes
 $c=\frac{3(3\nu-1)(4\nu-1)}{2\nu(\nu+1)}.$ (90)
-We observe that the fermionzation of each theory in (89) denoted as $\mathcal{F}\mathcal{T}_{2}^{(3,\nu)}$ has at most four distinct Neveu–Schwarz characters and equal number of Ramond characters.
-In general, we find the four Neveu–Schwarz weights are
+We observe that the fermionization of each theory in (89), denoted as $\mathcal{F}\mathcal{T}_{2}^{(3,\nu)}$, has at most four distinct Neveu–Schwarz characters and an equal number of Ramond characters.
+In general, we find that the four Neveu–Schwarz conformal weights are
 $0,\frac{4\nu-1}{4(\nu+1)},\frac{2\nu-1}{2\nu},\frac{7\nu^{2}-7\nu+6}{4\nu(\nu+1)},$ (91)
-while the four Ramond weights are
+while the four Ramond conformal weights are
 $\frac{(3\nu-1)(4\nu-1)}{16\nu(\nu+1)},\frac{3(4\nu-1)}{16\nu},\frac{(3\nu-1)(8\nu+3)}{16\nu(\nu+1)},\frac{3(8\nu-5)}{16\nu}.$ (92)
 Notice the first Ramond weight has exponent 0 which indicates the existence of the constant $\widetilde{\rm R}$ character.
-One can also easily recover 13 bosonic conformal weights and check that the 13 conformal weights always yield a 13th order MLDE with index $l=36$.
-For theories with less than four distinct NS characters, the fermionic weights take value in the above list.
-We further find a one-parameter family of holomorphic fermionic MLDE of degree four for the Neveu–Schwarz characters of $\mathcal{F}\mathcal{T}_{2}^{(3,\nu)}$ theories.
+One can also easily recover 13 bosonic conformal weights and verify that the 13 conformal weights always yield a 13th order MLDE with index $l=36$.
+For theories with less than four distinct NS characters, the fermionic weights take values in the list above.
+We further find a one-parameter family of holomorphic fermionic MLDE of degree four for the Neveu–Schwarz characters of the theories $\mathcal{F}\mathcal{T}_{2}^{(3,\nu)}$.
 To present the fermionic MLDE, we first recall the Jacobi theta functions
@@
 $+\lambda_{5}(\theta_{4}^{12}-\theta_{2}^{12})\big)D+\lambda_{6}(\theta_{2}^{16}+\theta_{4}^{16})+\lambda_{7}(\theta_{4}^{12}\theta_{2}^{4}+\theta_{2}^{12}\theta_{4}^{4})+\lambda_{8}\theta_{2}^{8}\theta_{4}^{8}\Big]\chi^{\rm NS}=0.$
-In the meantime, the Ramond characters should satisfy the following dual MLDE
+In the meantime, the Ramond characters satisfy the following dual MLDE obtained from modular transformation
 $\Big[D^{4}$ $-\lambda_{1}(\theta_{3}^{4}+\theta_{4}^{4})D^{3}+\big(\lambda_{2}(\theta_{4}^{8}+\theta_{3}^{8})-\lambda_{3}\theta_{4}^{4}\theta_{3}^{4}\big)D^{2}+\big(\lambda_{4}(\theta_{3}^{8}\theta_{4}^{4}+\theta_{4}^{8}\theta_{3}^{4})$ (95)
 - $\,\lambda_{5}(\theta_{3}^{12}+\theta_{4}^{12})\big)D+\lambda_{6}(\theta_{4}^{16}+\theta_{3}^{16})-\lambda_{7}(\theta_{3}^{12}\theta_{4}^{4}+\theta_{4}^{12}\theta_{3}^{4})+\lambda_{8}\theta_{4}^{8}\theta_{3}^{8}\Big]\chi^{\rm R}=0.$
-The eight fermionic weights (91) and (92) along with the no free fermion condition allow us to solve the unfixed coefficients $\lambda_{i}$ are rational functions of $\nu$.
-In the end, we find the one-parameter family as
+The eight fermionic weights (91) and (92) along with the no free fermion condition allow us to solve the unfixed coefficients $\lambda_{i}$ as rational functions of $\nu$.
+In the end, we find the one-parameter family given by
 $\lambda_{1}=\frac{\nu^{2}+3\nu-3}{4\nu(\nu+1)},\ \ \lambda_{2}=-\frac{856\nu^{4}-1654\nu^{3}+2827\nu^{2}-2610\nu+1053}{1152\nu^{2}(\nu+1)^{2}},$ (96)
@@
 $\lambda_{7}=\frac{3(3\nu-1)(4\nu-1)\left(32\nu^{6}-740\nu^{5}+1740\nu^{4}-2133\nu^{3}+1121\nu^{2}-207\nu+27\right)}{16384\nu^{4}(\nu+1)^{4}},$
-and $\lambda_{8}=2\lambda_{7}-2\lambda_{6}$.
-The last equation comes from the supersymmetry condition that one Ramond weight has zero exponent.
-We checked the Neveu–Schwarz and Ramond characters of $\mathcal{F}\mathcal{T}_{2}^{(3,\nu)}$ theories indeed satisfy the above fermionic MLDEs.
-This one-parameter family of fermionic MLDE suggests that all Neveu–Schwarz and Ramond characters of $\mathcal{F}\mathcal{T}_{2}^{(3,\nu)}$ theory can be solved as functions of $\nu$ in a uniform way.
+$\lambda_{8}=2\lambda_{7}-2\lambda_{6}.$
+The last equation comes from the supersymmetry condition that one Ramond conformal weight has zero exponent.
+We have explicitly checked that the Neveu–Schwarz and Ramond characters of the theories $\mathcal{F}\mathcal{T}_{2}^{(3,\nu)}$ indeed satisfy the above fermionic MLDEs.
+This one-parameter family of fermionic MLDE suggests that all Neveu–Schwarz and Ramond characters of the theories $\mathcal{F}\mathcal{T}_{2}^{(3,\nu)}$ can be solved uniformly as functions of $\nu$.
 4.2 Coset constructions
-At level two, there no longer exists a universal coset construction covering the entire magic triangle, in contrast to the level-one case.
-Nevertheless, the magic coset relation (48) continues to provide strong guidance and motivates us to search for new uniform coset constructions involving the Cvitanović–Deligne, sub and Severi exceptional series at level two.
-First we find the following universal coset construction between Cvitanović–Deligne and sub exceptional series at level two:
+In contrast to the level-one case, there no longer exists a universal coset construction covering the entire magic triangle at level two.
+Nevertheless, the magic coset relation (48) continues to provide strong guidance and motivates us to search for new uniform coset constructions involving the Cvitanović–Deligne, subexceptional, and Severi exceptional series at level two.
+We first find the following universal coset construction between the Cvitanović–Deligne and subexceptional series at level two:
 $\frac{\mathcal{T}^{(5,\nu)}_{2}}{\mathcal{T}^{(3,\nu)}_{2}\times\mathcal{T}^{(5,1/3)}_{2}}=\frac{(\mathfrak{g}_{\rm CD})_{2}}{(\mathfrak{g}_{\rm sub})_{2}\times(A_{1})_{2}}=\mathrm{Vir}_{n+1,n},\qquad n=\left\{\begin{array}[]{ll}3+\frac{4}{\nu-1},&\nu>1\\ \frac{4}{1-\nu}-4,&\nu<1.\end{array}\right$.
@@
 $\frac{(G_{2})_{2}}{(A_{1})_{6}\times(A_{1})_{2}}=\mathrm{Vir}_{9,8}.$ (100)
-Each of the four characters of $(G_{2})_{2}$ decomposes to the summation of 12 terms of $(A_{1})_{6}\times(A_{1})_{2}\times\mathrm{Vir}_{9,8}$.
-For instance, denote $\mathrm{Vir}_{9,8}$ as $M$, then for the vacuum we have
+Each of the four characters of $(G_{2})_{2}$ decomposes into the sum of 12 terms of $(A_{1})_{6}\times(A_{1})_{2}\times\mathrm{Vir}_{9,8}$.
+For instance, denote $\mathrm{Vir}_{9,8}$ as $M$.
+For the vacuum character, we have
 $\chi^{(G_{2})_{2}}_{0}=$ $\ \chi^{(A_{1})_{2}}_{0}\chi^{(A_{1})_{6}}_{0}\chi^{M}_{0}+\chi^{(A_{1})_{2}}_{\frac{3}{16}}\chi^{(A_{1})_{6}}_{\frac{15}{32}}\chi^{M}_{\frac{11}{32}}+\chi^{(A_{1})_{2}}_{0}\chi^{(A_{1})_{6}}_{\frac{3}{4}}\chi^{M}_{\frac{5}{4}}$ (101)
@@
 The twelve terms are ordered according to increasing total conformal weight.
-We further find the following universal coset construction between the sub and Severi exceptional series at level 2:
+We further find the following universal coset construction between the subexceptional and Severi exceptional series at level 2:
 $\frac{\mathcal{T}^{(3,\nu)}_{2}}{\mathcal{T}^{(2,\nu)}_{2}\times\mathcal{T}^{(3,1/2)}_{2}}=\frac{(\mathfrak{g}_{\rm sub})_{2}}{(\mathfrak{g}_{\rm Severi})_{2}\times(U_{1})_{6}}=\mathcal{B}\mathrm{Vir}_{n+2,n}^{N=1},\qquad n=\left\{\begin{array}[]{ll}4+\frac{4}{\nu-1},&\nu>1,\\ \frac{4}{1-\nu}-6,&\nu<1.\end{array}\right$.
@@
 Here again $\nu=\frac{1}{2},\frac{2}{3},\frac{3}{2},2,3,5$.
-The $\mathcal{B}\mathrm{Vir}_{n+2,n}^{N=1}$ is the bosonization $\mathcal{B}$ of the unitary $N=1$ supersymmetric minimal model $\mathrm{Vir}_{n+2,n}^{N=1}$.
-For example, taking $\nu=3$ gives
+The theory $\mathcal{B}\mathrm{Vir}_{n+2,n}^{N=1}$ denotes the bosonization $\mathcal{B}$ of the unitary $N=1$ supersymmetric minimal model $\mathrm{Vir}_{n+2,n}^{N=1}$.
+As an example for $\nu=3$ the coset (104) gives
 $\frac{(E_{7})_{2}}{(E_{6})_{2}\times(U_{1})_{6}}=\mathcal{B}\mathrm{Vir}_{7,5}^{N=1}$ (105)
-It is easy to find $\mathcal{B}\mathrm{Vir}_{7,5}^{N=1}$ has in total $18$ bosonic primary fields with conformal weights
+It is straightforward to verify that find $\mathcal{B}\mathrm{Vir}_{7,5}^{N=1}$ has in total $18$ bosonic primary fields with conformal weights
 $\left\{0,\frac{3}{70},\frac{29}{560},\frac{9}{112},\frac{4}{35},\frac{3}{14},\frac{27}{80},\frac{269}{560},\frac{19}{35},\frac{43}{70},\frac{73}{112},\frac{5}{7},\frac{9}{10},\frac{8}{7},\frac{7}{5},\frac{3}{2},\frac{23}{14},\frac{31}{16}\right\}.$
-Denote $\mathcal{B}\mathrm{Vir}_{7,5}^{N=1}$ as $\mathcal{B}$, then for the vacuum, we find the character character relation
+Denote $\mathcal{B}\mathrm{Vir}_{7,5}^{N=1}$ as $\mathcal{B}$, the vacuum character satisfies the following relation:
 $\chi^{(E_{7})_{2}}_{0}=$ $\ \chi^{(E_{6})_{2}}_{0}\chi^{(U_{1})_{6}}_{0}\chi^{\mathcal{B}}_{0}+2\chi^{(E_{6})_{2}}_{\frac{13}{21}}\chi^{(U_{1})_{6}}_{\frac{1}{6}}\chi^{\mathcal{B}}_{\frac{3}{14}}+\chi^{(E_{6})_{2}}_{\frac{6}{7}}\chi^{(U_{1})_{6}}_{0}\chi^{\mathcal{B}}_{\frac{8}{7}}$ (106)
@@
 $+\chi^{(E_{6})_{2}}_{0}\chi^{(U_{1})_{6}}_{\frac{3}{2}}\chi^{\mathcal{B}}_{\frac{3}{2}}+\chi^{(E_{6})_{2}}_{\frac{9}{7}}\chi^{(U_{1})_{6}}_{\frac{3}{2}}\chi^{\mathcal{B}}_{\frac{3}{14}}+\chi^{(E_{6})_{2}}_{\frac{6}{7}}\chi^{(U_{1})_{6}}_{\frac{3}{2}}\chi^{\mathcal{B}}_{\frac{23}{14}}.$
-We check the degeneracy matches and it is straightforward to upgrade the character relation to module decomposition.
-Furthermore, there exists the following universal coset Cheng et al. (2020) for Cvitanović–Deligne exceptional series resembling the Maverick coset
+We have checked that the degeneracies match, and it is straightforward to upgrade the character relation to module decomposition.
+Furthermore, there exists the following universal coset Cheng et al. (2020) for Cvitanović–Deligne exceptional series resembling the Maverick coset Dunbar and Joshi (1993)
 $\frac{\mathcal{T}^{(5,\nu)}_{1}\times\mathcal{T}^{(5,\nu)}_{1}}{\mathcal{T}^{(5,\nu)}_{2}}=\frac{(\mathfrak{g}_{\rm CD})_{1}\times(\mathfrak{g}_{\rm CD})_{1}}{(\mathfrak{g}_{\rm CD})_{2}}=\mathrm{Vir}_{n+1,n},\qquad n=\left\{\begin{array}[]{ll}2+\frac{4}{\nu-1},&\nu>1\\ \frac{4}{1-\nu}-3,&\nu<1.\end{array}\right$.
 (109)
-Here $\nu$ takes value in $1/3,1/2,2/3,3/2,2,3,5$.
-For $\nu=1/4,4$, it still holds if we scale $n$ and $n+1$ such that they are coprime integers and take the proper non-diagonal modular invariant.
-For $\nu=4$, the non-diagonal modular invariant was identified in Lee et al.
+Here $\nu=\frac{1}{2},\frac{2}{3},\frac{3}{2},2,3,5$.
+For $\nu=\frac{1}{4},4$, the relation continues to hold provided one scales $n$ and $n+1$ so that they are coprime integers and choose the appropriate non-diagonal modular invariant.
+For $\nu=4$, such a non-diagonal modular invariant was identified in Lee et al.
 (2024a).
-For $\nu=1/4$, the coset was recently proved in Feng and Jiang (2025).
-We newly find a similar coset construction for the subexceptional series:
-$\frac{(\mathfrak{g}_{\rm sub})_{1}\times(\mathfrak{g}_{\rm sub})_{1}}{(\mathfrak{g}_{\rm sub})_{2}}=\mathcal{B}\mathrm{Vir}_{n+2,n}^{N=1},\qquad n=\left\{\begin{array}[]{ll}2+\frac{4}{\nu-1},&\nu>1,\\ \frac{4}{1-\nu}-4,&\nu<1.\end{array}\right$.
+For $\nu=\frac{1}{4}$, the corresponding coset was recently proved in Feng and Jiang (2025), with $\mathfrak{g}^{(5,1/2)}$ as $osp(1|2)$.
+We further find a similar coset construction for the subexceptional series:
+$\frac{\mathcal{T}^{(3,\nu)}_{1}\times\mathcal{T}^{(3,\nu)}_{1}}{\mathcal{T}^{(3,\nu)}_{2}}=\frac{(\mathfrak{g}_{\rm sub})_{1}\times(\mathfrak{g}_{\rm sub})_{1}}{(\mathfrak{g}_{\rm sub})_{2}}=\mathcal{B}\mathrm{Vir}_{n+2,n}^{N=1},\qquad n=\left\{\begin{array}[]{ll}2+\frac{4}{\nu-1},&\nu>1,\\ \frac{4}{1-\nu}-4,&\nu<1.\end{array}\right$.
 (112)
-We have a final remark for the intermediate exceptional series at level 2. For $\nu=2,3,5$ and $n=3+\frac{4}{\nu-1}$, the following cosets were found in (Sun, 2024, Equation (1.9),(1.10)):
-$\frac{\mathcal{T}^{(5,\nu)}_{2}}{\mathcal{T}^{(4,\nu)}_{2}}$ $=\frac{(\mathfrak{g}_{\rm CD})_{2}}{(\mathfrak{g}_{I})_{2}}=(\mathrm{Vir}_{3n+1,n})_{\rm eff},$ (113)
-$\frac{\mathcal{T}^{(4,\nu)}_{2}}{\mathcal{T}^{(3,\nu)}_{2}}$ $=\frac{(\mathfrak{g}_{I})_{2}}{(\mathfrak{g}_{\rm sub})_{2}}=(\mathcal{B}\mathrm{Vir}^{N=1}_{3n+1,n+1})_{\rm eff}.$ (114)
+Finally, we make a remark concerning the intermediate exceptional series at level 2. For $\nu=2,3,5$ and $n=3+\frac{4}{\nu-1}$, the following cosets were found in (Sun, 2024, Equation (1.9),(1.10)):
+$\frac{\mathcal{T}^{(5,\nu)}_{2}}{\mathcal{T}^{(4,\nu)}_{2}}$ $=\frac{(\mathfrak{g}_{\rm CD})_{2}}{(\mathfrak{g}_{I})_{2}}=(\mathrm{Vir}_{3n+1,n})^{\rm eff},$ (113)
+$\frac{\mathcal{T}^{(4,\nu)}_{2}}{\mathcal{T}^{(3,\nu)}_{2}}$ $=\frac{(\mathfrak{g}_{I})_{2}}{(\mathfrak{g}_{\rm sub})_{2}}=(\mathcal{B}\mathrm{Vir}^{N=1}_{3n+1,n+1})^{\rm eff}.$ (114)
 5 Magic triangle at arbitrary positive integer levels
-Denote the RCFT associated with the triangle element $(\mu,\nu)$ at level $k$ as $\mathcal{T}^{(\mu,\nu)}_{k}$.
-For the 16 elements of the magic triangle that are semisimple Lie algebras $\mathfrak{g}$, the corresponding RCFT with arbitrary positive integer level $k$ are just their WZW models $(\mathfrak{g})_{k}$, except for ${\mathfrak{g}}^{(3,2/3)}=A_{1}$, ${\mathfrak{g}}^{(3/2,3/2)}=A_{1}$ and ${\mathfrak{g}}^{(2,3/2)}=A_{2}$ for which the levels must be rescaled by $3$, 88, and $2$, respectively.
+We denote the RCFT associated with the triangle element $(\mu,\nu)$ at level $k$ as $\mathcal{T}^{(\mu,\nu)}_{k}$.
+For the 16 elements of the magic triangle that are semisimple Lie algebras $\mathfrak{g}$, the corresponding RCFT with arbitrary positive integer level $k$ are just their WZW models $(\mathfrak{g})_{k}$, except for ${\mathfrak{g}}^{(3,2/3)}=A_{1}$, ${\mathfrak{g}}^{(3/2,3/2)}=A_{1}$ and ${\mathfrak{g}}^{(2,3/2)}=A_{2}$ for which the levels must be rescaled by factors $3$, 88, and $2$, respectively.
 For the remaining 14 elements, a case-by-case analysis is required.
@@
 (2022).
-  • $\mathcal{T}^{(1/3,4)}_{k}$ is the effective minimal model $\mathrm{Vir}_{3k+2,3}^{\rm eff}$. This has been realized in Lee et al. (2024b).
-  • $\mathcal{T}^{(1/2,3)}_{k}$ is the compact boson $U(1)_{3k}$ with $c=1$.
+  • $\mathcal{T}^{(1/3,4)}_{k}$ is the effective minimal model $\mathrm{Vir}_{3k+2,3}^{\rm eff}$. This was realized in Lee et al. (2024b).
+  • $\mathcal{T}^{(1/2,3)}_{k}$ is the compact boson $(U_{1})_{3k}$ with $c=1$. The central charge is due to $d_{\mathfrak{g}}=1$ and $h^{\vee}=0$.
   • $\mathcal{T}^{(2/3,2)}_{k}$ and $\mathcal{T}^{(1,3/2)}_{k}$ are trivial for $k>1$ due to $d=0$ and $h^{\vee}=-1$.
-  • $\mathcal{T}^{(1,2)}_{k}$ is a non-diagonal modular invariant of $U(1)_{2k}\times U(1)_{6k}$. The vacuum character is given by
-    $\chi^{U(1)_{2k}}_{0}\chi^{U(1)_{6k}}_{0}+\chi^{U(1)_{2k}}_{\frac{k}{2}}\chi^{U(1)_{6k}}_{\frac{3k}{2}}.$ (115)
+  • $\mathcal{T}^{(1,2)}_{k}$ is a non-diagonal modular invariant of $(U_{1})_{2k}\times(U_{1})_{6k}$ with $c=2$.
+The central charge is due to $d_{\mathfrak{g}}=2$ and $h^{\vee}=0$.
+The vacuum character is given by
+    $\chi^{(U_{1})_{2k}}_{0}\chi^{(U_{1})_{6k}}_{0}+\chi^{(U_{1})_{2k}}_{\frac{k}{2}}\chi^{(U_{1})_{6k}}_{\frac{3k}{2}}.$ (115)
   • $\mathcal{T}^{(3/2,3/2)}_{k}$ is the $D$-type non-diagonal modular invariant of $(A_{1})_{8k}$.
-  • $\mathcal{T}^{(4,5)}_{k}$, $\mathcal{T}^{(4,3)}_{k}$, $\mathcal{T}^{(4,2)}_{k}$ exist up to $k=5,3,2$ respectively.
-These theories are equivalent to the minimal $\mathcal{W}$-algebras of E-type with negative integer levels, which have recently been studied in Sun (2024); Arakawa et al.
+  • The theories $\mathcal{T}^{(4,5)}_{k}$, $\mathcal{T}^{(4,3)}_{k}$, and $\mathcal{T}^{(4,2)}_{k}$ exist up to $k=5,3,2$ respectively.
+These theories are equivalent to the minimal $\mathcal{W}$-algebras of E-type with negative integer levels, which have recently been studied in Arakawa and Moreau (2018); Sun (2024); Arakawa et al.
 (2024); Kac and Wakimoto (2025).
-Other $\mathcal{T}^{(4,\nu)}_{k}$ theories for $k>1$ should not exist.
+Other $\mathcal{T}^{(4,\nu)}_{k}$ theories for $k>1$ are not expected to exist.
 6 Concluding remarks
 In this work, we have extended the magic triangle and identified the two-dimensional conformal field theories associated with its elements at arbitrary positive integer levels.
-In particular, for level-one theories we uncover a number of remarkable and highly structured properties, including modular linear differential equations, the magic coset construction, five atomic theories, magic bilinear relations, and matrix-valued modularity.
+In particular, for level-one theories we uncover a number of remarkable properties, including modular linear differential equations, the magic coset, five atomic theories, magic bilinear relations, the dimensions and degeneracies of primary operators, and matrix-valued modularity including $S$-matrices.
 For level-two theories, we focus on the subexceptional series and found they are all supersymmetric and their fermionic characters satisfy a uniform fermionic MLDE.
@@
 We have not explored much of the dimension of irreducible representations for the magic triangle in this work.
-The magic coset we have uncovered suggest that the representations of all algebras appearing in the magic triangle are intricately entangled.
-It would be interesting to find some simple universal formulas for the dimension and dual Coxeter number of these representations, besides the adjoint representation.
-Finally, although our discussion has focused on WZW models at positive integer levels, theories at fractional and negative levels are also of significant interest.
+The magic coset we have uncovered suggest that the representations of all algebras appearing in the magic triangle are intricately interconnected.
+It would therefore be interesting to find some simple universal formulas for the dimensions of these representations, besides the adjoint representation and some simpler representations appearing as those for non-vacuum primary operators.
+Finally, although our discussion has focused on WZW models at positive integer levels, theories at fractional and negative levels are also of considerable interest.
 For instance, it is well known that the Cvitanović–Deligne exceptional series at level $h^{\vee}/6-1$ gives rise to a celebrated family of quasi-lisse vertex operator algebras, whose vacuum characters satisfy a one-parameter family of second-order MLDEs Beem and Rastelli (2018); Arakawa and Kawasetsu (2018).
@@
 KL is supported in part by the BIMSA start-up fund and the Beijing Natural Science Foundation International Scientist Project No.IS25024.
-KS is supported by NSF No.1250010384.
+KS is supported by China NSF grant No.1250010384.
 Appendix A Explicit data of the extended magic triangle
 We collect the dual Coxeter number of dimension of each element of the extended magic triangle in Table 5.
-We collect the dimensions of $R_{1},R_{2},R_{3}$ for all theories at level one in Table 6.
-We collect the central charge, nonzero conformal weights and degeneracies for all theories at level one in Table 7.
-Due to the symmetry between $\mu$ and $\nu$, we only show the $\mu\leq\nu$ part of the triangle.
+We collect the dimensions of $R_{1},R_{2},R_{3}$ for all level-one theories in Table 6.
+The central charge, nonzero conformal weights and degeneracies for all level-one theories are summarized in Table 7.
+Owing to the symmetry exchanging $\mu$ and $\nu$, we display the $\mu\leq\nu$ portion of the triangle.
 $\{-1,0\}$ $\left\{\frac{1}{4},3\right\}$ $3/2$
@@
 Appendix B S-matrices
-In this section, we collect the $S$-matrices for the two characters $\chi_{0},\chi_{1}$ of the Mathur–Mukhi–Sen series, i.e.
-the extended Cvitanović–Deligne exceptional series at level one, reviewed in Section 2.3.
-The $2\times 2$ $S$-matrix $S_{(\nu)}$ for the two characters of $\mathcal{T}^{(\nu,5)}$ which leaves the partition function (15) invariant satisfies $S_{(\nu)}=S_{(1/\nu)}.$ The $S$-matrices for $\nu=1/4,1/3,1/2,2/3,1$ of ${\rm Vir}_{5,2}^{\rm eff},(A_{1})_{1},(A_{2})_{1},(G_{2})_{1},(D_{4})_{1}$ theories are, respectively,
-$S_{(4)}=\frac{\varphi^{1/2}}{5^{1/4}}\left(\begin{array}[]{cc}1&\varphi^{-1}\\ \varphi^{-1}&-1\end{array}\right),\quad S_{(3)}=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cr}1&1\\ 1&-1\end{array}\right),\quad S_{(2)}=\frac{1}{\sqrt{3}}\left(\begin{array}[]{cr}1&2\\ 1&-1\end{array}\right),$ (122)
-$S_{(3/2)}=\frac{1}{5^{1/4}\varphi^{1/2}}\left(\begin{array}[]{cc}1&\varphi\\ \varphi&-1\end{array}\right),\quad S_{(1)}=\frac{1}{2}\left(\begin{array}[]{cr}1&3\\ 1&-1\end{array}\right),$ (127)
-where $\varphi=(\sqrt{5}+1)/2$ is the Golden ratio. The partition function $Z=|\chi_{0}|^{2}+\mathcal{N}|\chi_{1}|^{2}$ is invariant under these $S$ matrices. Note that
+In this section, we collect the $S$-matrices for the two characters $\chi_{0},\chi_{1}$ of the Mathur–Mukhi–Sen series, namely the extended Cvitanović–Deligne exceptional series at level one, as reviewed in Section 2.3.
+The $2\times 2$ $S$-matrix $S_{(\nu)}$ for the two characters of $\mathcal{T}^{(\nu,5)}$, which leaves the partition function (15) invariant, satisfies $S_{(\nu)}=S_{(1/\nu)}.$ The $S$-matrices for $\nu=1/4,1/3,1/2,2/3,1$, corresponding respectively to the theories ${\rm Vir}_{5,2}^{\rm eff}$, $(A_{1})_{1}$, $(A_{2})_{1}$, $(G_{2})_{1}$, $(D_{4})_{1}$, are given by
+$S_{(1/4)}=\frac{\varphi^{1/2}}{5^{1/4}}\left(\begin{array}[]{cc}1&\varphi^{-1}\\ \varphi^{-1}&-1\end{array}\right),\quad S_{(1/3)}=\frac{1}{\sqrt{2}}\left(\begin{array}[]{cr}1&1\\ 1&-1\end{array}\right),\quad S_{(1/2)}=\frac{1}{\sqrt{3}}\left(\begin{array}[]{cr}1&2\\ 1&-1\end{array}\right),$ (122)
+$S_{(2/3)}=\frac{1}{5^{1/4}\varphi^{1/2}}\left(\begin{array}[]{cc}1&\varphi\\ \varphi&-1\end{array}\right),\quad S_{(1)}=\frac{1}{2}\left(\begin{array}[]{cr}1&3\\ 1&-1\end{array}\right),$ (127)
+where $\varphi=(\sqrt{5}+1)/2$ is the golden ratio.
+The modular invariant partition function is $Z=|\chi_{0}|^{2}+\mathcal{N}|\chi_{1}|^{2}$ and $\mathcal{N}$ is the degeneracy of the non-vacuum primary fields.
+Note that
 $S_{(\nu)}^{T}\cdot\left(\begin{array}[]{cc}1&0\\ 0&\mathcal{N}_{(\nu)}\end{array}\right)\cdot S_{(\nu)}=\left(\begin{array}[]{cc}1&0\\ 0&\mathcal{N}_{(\nu)}.\end{array}\right)$ (132)
 In addition for arbitrary $\nu$ we have $\det S_{(\nu)}=-1.$
-For the cases with degeneracy, i.e. $\nu=1/2,1,2$, we can have extended $S$-matrices for $\nu=1/2,1$ as
-$\frac{1}{\sqrt{3}}\left(\begin{array}[]{ccc}1&1&1\\ 1&w&w^{2}\\ 1&w^{2}&w\end{array}\right),\quad\frac{1}{2}\left(\begin{array}[]{crrr}1&1&1&1\\ 1&1&-1&-1\\ 1&-1&1&-1\\ 1&-1&-1&1\end{array}\right),$ (140)
+For the cases with degeneracy, namely $\nu=1/2,1,2$, we can have extended $S$-matrices, which are given by
+$S_{(1/2)}=\frac{1}{\sqrt{3}}\left(\begin{array}[]{ccc}1&1&1\\ 1&w&w^{2}\\ 1&w^{2}&w\end{array}\right),\quad\ S_{(1)}=\frac{1}{2}\left(\begin{array}[]{crrr}1&1&1&1\\ 1&1&-1&-1\\ 1&-1&1&-1\\ 1&-1&-1&1\end{array}\right),$ (140)
 where $w=e^{2\pi i/3}$.
@@
 Appendix C Spurious solutions of MLDE
-For each $\mu$, besides the two characters of $\mathcal{T}^{(5,\mu)}$, there are two spurious solutions of the fourth order MLDE (41) with Fourier coefficients with unbounded denominators.
-For example, for $\mu=1/4$, the two genuine solutions are the Lee–Yang characters $\chi_{0},\chi_{1/5}$, while the two spurious solutions are
+For each $\mu$, in addition to the two genuine characters of $\mathcal{T}^{(5,\mu)}$, there exist two spurious solutions of the fourth order MLDE (41) whose q-expansion have coefficients with unbounded denominators.
+As an illustrative example, for $\mu=1/4$, the two genuine solutions are the Lee–Yang characters $\chi_{0},\chi_{1/5}$, while the two spurious solutions are given by
 $f_{\frac{1}{30}}=$ $\ q^{\frac{1}{60}}\Big(1-\frac{2q}{5}+\frac{q^{2}}{11}+\frac{26q^{3}}{85}+\frac{434q^{4}}{1265}+\frac{9824q^{5}}{27115}+\dots\Big),$
 $f_{\frac{5}{6}}=$ $\ q^{\frac{49}{60}}\Big(1+\frac{38q}{33}+\frac{371q^{2}}{561}+\frac{22558q^{3}}{12903}+\frac{383219q^{4}}{374187}+\dots\Big).$
-Although these two solutions are not characters themselves, we observe that they are closely related to the characters of the $(E_{7+1/2})_{1}$ theory:
+Although these two solutions are not themselves RCFT characters, we observe that they are closely related to the characters of the $(E_{7+1/2})_{1}$ theory via the relations
 $-180\Delta_{\frac{1}{4}}f_{\frac{1}{30}}=$ $\,\eta^{8}\chi_{0}^{(E_{7+1/2})_{1}},$
 $108\Delta_{\frac{1}{4}}f_{\frac{5}{6}}=$ $\,\eta^{8}\chi_{\frac{4}{5}}^{(E_{7+1/2})_{1}}.$
-These relations are easy to prove since $\Delta_{4}$ annihilates both sides of the equations.
-If we define $\chi_{\frac{1}{30}}=R_{1}f_{\frac{1}{30}}$ and $\chi_{\frac{5}{6}}=R_{3}f_{\frac{5}{6}}$, by properly choosing the prefactors such that $R_{3}/R_{1}=3/5$, we find a bilinear relation
+Here $\Delta_{\mu}$ is the second order differential operator defined in (14).
+These relations are straightforward to prove, since $\Delta_{4}$ annihilates both sides of the equations.
+If we define $\chi_{\frac{1}{30}}=R_{1}f_{\frac{1}{30}}$ and $\chi_{\frac{5}{6}}=R_{3}f_{\frac{5}{6}}$, and choose the prefactors such that $R_{3}/R_{1}=3/5$, we obtain the bilinear relation
 $\chi_{0}\chi_{\frac{1}{30}}-\chi_{\frac{1}{5}}\chi_{\frac{5}{6}}=R_{1}.$ (146)
-similar to the magic bilinear relation (46). This choice also makes the relation (47) hold. For other $\mu\neq 5$, we obtain similar relations.
-The level one $E_{8}$ theory has only one character $\chi_{0}=j^{1/3}(\tau)$ where $j(\tau)$ is the Klein $j$-function. The vacuum character satisfies the first order MLDE,
+similar to the magic bilinear relation (46). This choice also ensures the relation (47) holds. For other $\mu\neq 5$, similar relations can be derived.
+The level one $(E_{8})_{1}$ theory has a single character $\chi_{0}=j^{1/3}(\tau)$ where $j(\tau)$ is the Klein $j$-invarant. The vacuum character satisfies the first order meromorphic MLDE,
 $\Big(D+\frac{E_{6}}{3E_{4}}\Big)j^{1/3}=0.$ (147)
-The second order MLDE (10) for $E_{8}$ theory gets factorized as
+Accordinngly, the second-order MLDE (10) for the $E_{8}$ theory factorizes as
 $D^{2}-\frac{E_{4}}{6}=\Big(D-\frac{E_{6}}{3E_{4}}\Big)\Big(D+\frac{E_{6}}{3E_{4}}\Big).$ (148)
-For $\mu,\nu=5$, the 4th order MLDE (41) becomes
-$\left[D^{4}-\frac{1}{6}E_{4}D^{2}+\frac{1}{9}E_{6}D-\frac{1}{36}E_{4}^{2}\right]\chi=0.$ (149)
-It factorizes to $D^{2}(D^{2}-\frac{1}{6}E_{4})]\chi=0$. Beside $\chi_{0}=j^{1/3}$ the vacuum character of $(E_{8})_{1}$, this MLDE has two more spurious solutions
+For $\mu,\nu=5$, the fourth order MLDE (41) becomes
+$\left[D^{4}-\frac{1}{6}E_{4}D^{2}+\frac{1}{9}E_{6}D-\frac{1}{36}E_{4}^{2}\right]\chi=\left[D^{2}(D^{2}-\frac{1}{6}E_{4})\right]\chi=0.$ (149)
+Besides $\chi_{0}=j^{1/3}$ the vacuum character of $(E_{8})_{1}$, this MLDE has two more spurious solutions
 $f_{\frac{2}{3}}=$ $\ q^{\frac{1}{3}}\Big(1+\frac{712q}{25}+\frac{75436q^{2}}{275}+\frac{95654016q^{3}}{51425}+\dots\Big),$
 $f_{\frac{5}{6}}=$ $\ q^{\frac{1}{2}}\Big(1+\frac{228q}{11}+\frac{34938q^{2}}{187}+\frac{5163352q^{3}}{4301}+\dots\Big).$
-These are not characters, but satisfy the following simple relations
+These solutions are not characters, but satisfy the following simple differential equations
 $-9\Big(D^{2}-\frac{1}{6}E_{4}\Big)f_{\frac{2}{3}}=$ $\ \eta^{8},$
 $\frac{6}{5}\Big(D+\frac{E_{6}}{3E_{4}}\Big)f_{\frac{5}{6}}=$ $\ \eta^{4}j^{-1/3}.$
-We find a bilinear relation
+We find a bilinear relation for the three solutions
 $\chi_{0}f_{\frac{2}{3}}-\frac{2^{8}\cdot 3^{3}}{5^{2}}f_{\frac{5}{6}}^{2}=1,$ (150)